### v2

# SemGauss-SLAM: Dense Semantic Gaussian Splatting SLAM

Abstract
We propose SemGauss-SLAM, the first semantic SLAM system utilizing 3D Gaussian representation, that enables accurate 3D semantic mapping, robust camera tracking, and high-quality rendering in real-time. In this system, we incorporate semantic feature embedding into 3D Gaussian representation, which effectively encodes semantic information within the spatial layout of the environment for precise semantic scene representation. Furthermore, we propose feature-level loss for updating 3D Gaussian representation, enabling higher-level guidance for 3D Gaussian optimization. In addition, to reduce cumulative drift and improve reconstruction accuracy, we introduce semantic-informed bundle adjustment leveraging semantic associations for joint optimization of 3D Gaussian representation and camera poses, leading to more robust tracking and consistent mapping. Our SemGauss-SLAM method demonstrates superior performance over existing dense semantic SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in novel-view semantic synthesis and 3D semantic mapping. The source code will be released upon acceptance.
Keywords:
Dense Semantic SLAM Semantic-informed Bundle Adjustment 3D Gaussian Semantic Representation 3D Reconstruction
1 Introduction
Dense semantic Simultaneous Localization and Mapping (SLAM) is a fundamental challenge for robotic systems [18, 15] and autonomous driving [1, 10]. It integrates semantic understanding of the environment into dense map reconstruction and performs pose estimation simultaneously. Traditional semantic SLAM has limitations including its inability to predict unknown areas and the demand for considerable map storage [11]. Subsequent semantic SLAM based on Neural Radiance Fields (NeRF) [16] methods [32, 9] have addressed these drawbacks, but suffers from inefficient per-pixel raycasting rendering.
Recently, a novel radiance field based on 3D Gaussian [8] has demonstrated remarkable capability in scene representation, enabling high-quality and efficient rendering via splatting. Following the advantages of 3D Gaussian representation, 3D Gaussian SLAM methods [27, 7, 31, 13] have been developed to achieve photo-realistic mapping. However, existing 3D Gaussian SLAM systems focus on visual mapping to obtain RGB maps, where color information alone is insufficient for downstream tasks such as navigation. Meanwhile, current semantic SLAM methods based on NeRF are prone to cumulative drift, leading to degraded SLAM accuracy. Therefore, developing a low-drift semantic SLAM system based on radiance field is essential and challenging.
Figure 1: Our SemGauss-SLAM incorporates semantic feature embedding into 3D Gaussian representation to perform dense semantic SLAM. This modeling strategy not only achieves accurate semantic mapping, but also enables high-precision semantic novel view synthesis compared with other radiance field-based semantic SLAM. We visualize 3D Gaussian blobs with semantic embedding, showing the spatial layout of semantic Gaussian representation. Moreover, semantic mapping is visualized using semantic feature embedding, showing 3D semantic modeling of the scene.
For radiance field-based semantic SLAM, there are two challenges: 1) Existing semantic SLAM methods struggle to achieve real-time 3D semantic mapping from 2D semantic information without predefined bounds, which is a fundamental requirement for semantic SLAM system. 2) As the progression of tracking, pose estimation is prone to cumulative drift, leading to a decrease in SLAM accuracy.
For the first challenge, SNI-SLAM [32] utilizes feature collaboration to achieve semantic mapping but requires specific scene bounds. Moreover, solely utilizing rendering results for semantic representation optimization without initial value is computationally inefficient. In this paper, we leverage the explicit structure of 3D Gaussian representation and propagate extracted 2D semantic features directly to 3D Gaussian as the initial value for achieving unbounded mapping and efficient 3D semantic scene optimization. This design enables faster transfer from 2D semantic information to 3D semantic map. In addition, we introduce 2D feature-level loss to guide scene representation optimization at a higher-level, thereby accelerating the convergence of 3D Gaussian optimization.
For the second challenge, NICE-SLAM [33] performs local bundle adjustment (BA) for joint optimization of pose and scene representation to reduce accumulated drift. However, it only utilizes appearance and geometry constraints, lacking adequate semantic constraints for semantic SLAM optimization. To address this challenge, we leverage semantic associations among co-visible frames and propose semantic-informed bundle adjustment for joint optimization of camera poses and 3D Gaussian representation. This design exploits the consistency of multi-view semantics for establishing constraints, which enables the reduction of cumulative drift in tracking and enhanced semantic mapping precision.
Overall, we provide the following contributions:
  • We present SemGauss-SLAM, the first 3D Gaussian semantic SLAM system, which can achieve accurate semantic mapping and photo-realistic reconstruction. We incorporate semantic feature embedding into 3D Gaussian for precisely constructing semantic maps. Moreover, feature-level loss is introduced to provide higher-level guidance for 3D Gaussian optimization.
  • We perform semantic-informed bundle adjustment by leveraging multi-view semantic constraints for joint optimization of camera poses and 3D Gaussian representation, achieving low-drift tracking and accurate semantic mapping.
  • We conduct extensive evaluations on two challenging datasets, Replica [21] and ScanNet [2], to demonstrate our method achieves state-of-the-art performance compared with existing radiance field-based SLAM in mapping, tracking, semantic segmentation and novel-view synthesis.
2 Related Work
Semantic SLAM can be divided into two main categories based on the form of scene representation: traditional semantic SLAM and neural implicit semantic SLAM. Traditional semantic SLAM employs explicit 3D representation such as surfels [15], mesh [18, 24, 5] and Truncated Signed Distance Fields (TSDF) [20, 14, 3]. Neural implicit semantic SLAM [4, 9, 32] utilizes implicit representation, such as feature grid [9] and feature plane [32].
Traditional Semantic SLAM.   SemanticFusion [15] uses surfel representation and employs Conditional Random Field (CRF) for updating class probability distribution incrementally. Fusion++ [14] performs object-level SLAM where each object is reconstructed within its own TSDF volume and segmented based on estimated foreground probability. Kimera [18] utilizes visual-inertial odometry for pose estimation and generates dense semantic mesh maps. However, these explicit scene modeling methods not only require high storage space but also fail to achieve high-fidelity and complete reconstruction.
Neural Implicit SLAM.   iMAP [23] first achieves real-time mapping and tracking utilizing a single MLP network for scene representation. To overcome oversmoothed scene reconstruction and improve scalability, NICE-SLAM [33] adopts hierarchical feature grid representation. Following this, several works [29, 6, 25, 19] introduce more efficient scene representation, such as hash-based feature grid and feature plane, to achieve more accurate SLAM performance.
For neural implicit semantic SLAM, DNS SLAM [9] utilizes 2D semantic priors and integrates multi-view geometry constraints for semantic reconstruction. SNI-SLAM [32] introduces feature collaboration and one-way correlation decoder for improved scene representation in semantic mapping. However, these methods require specific scene bounds for mapping and suffer from cumulative drift in pose estimation. In this paper, we leverage the explicit structure of 3D Gaussian for unbounded mapping, and perform semantic-informed bundle adjustment utilizing multi-frame semantic constraints for conducting low-drift, high-quality dense semantic SLAM.
3D Gaussian Splatting SLAM.   3D Gaussian [8] emerges as a promising 3D scene representation using a set of 3D Gaussians with position, anisotropic covariance, opacity, and color. This representation is capable of quick differential rendering through splatting and has a wide range of applications in dynamic scene modeling [12, 26, 28] and scene editing [34, 30].
Our main focus is on 3D Gaussian SLAM [27, 7, 13, 31]. These works emerge concurrently and all perform dense visual SLAM by leveraging scene geometry and appearance modeling capabilities of 3D Gaussian representation. SplaTAM [7] introduces silhouette-guided optimization to facilitate structured map expansion for dense mapping of visual SLAM. Gaussian Splatting SLAM [13] performs novel Gaussian insertion and pruning for monocular SLAM. However, the capability of 3D Gaussian representation extends well beyond appearance and geometry modeling of scene, as it can be augmented to perform semantic scene understanding. In this paper, we incorporate semantic feature embedding into 3D Gaussian for 3D semantic scene modeling to achieve high-precision dense semantic SLAM.
Figure 2: An overview of SemGauss-SLAM. Our method takes an RGB-D stream as input. RGB images are fed into feature extractor to obtain semantic features. These features are then categorized by pretrained classifier to attain semantic labels. Then, semantic features, semantic labels, along with the input RGB and depth data serve as supervision signals. In the meantime, semantic features and input RGB-D data propagate to 3D Gaussian blobs as initial properties of Gaussian representation. Rendered semantic feature, RGB, and depth are obtained from 3D Gaussian splatting, while rendered semantic label is attained by classifying rendered feature. Supervision and rendered information are utilized for loss construction to optimize camera poses and 3D Gaussian representation. During the SLAM process, we utilize semantic-informed bundle adjustment based on multi-frame constraints for joint optimization of poses and 3D Gaussian representation.
3 Method
SemGauss-SLAM is the first RGB-D dense semantic SLAM based on 3D Gaussian splatting. The overview of SemGauss-SLAM is shown in Fig. 2. Sec. 3.1 introduces semantic Gaussian representation for dense semantic mapping, as well as the process of 3D Gaussian-based semantic SLAM. Sec. 3.2 introduces the loss functions in the mapping and tracking process. Sec. 3.3 presents semantic-informed bundle adjustment to reduce accumulated drift in tracking and improve semantic mapping accuracy.
(a) 3D Gaussian Semantic Mapping
(b) Semantic-informed Bundle Adjustment
Figure 3: (a) 3D Gaussian semantic mapping process that utilizes 3D Gaussian blobs with semantic feature embedding for visualization. (b) Illustration of semantic-informed bundle adjustment.
3.1 3D Gaussian Semantic Mapping and Tracking
Semantic Gaussian Representation.   We utilize a set of Gaussians with specific properties for scene representation. To simplify scene representation for the SLAM process, we use isotropy Gaussians. Considering semantic representation, which has never been explored in other Gaussian SLAM, we introduce a new parameter, semantic feature embedding, to each Gaussian for semantic representation. To achieve real-time semantic mapping, it is crucial for 3D Gaussian representation, augmented by semantic feature embedding, to converge rapidly during the mapping optimization process. Therefore, instead of initializing these feature embeddings randomly, we propagate 2D semantic features extracted from images to 3D Gaussian as the initial values to achieve faster convergence of Gaussian semantic feature optimization. Moreover, integrating feature embedding into 3D Gaussian makes Gaussian semantic representation compact and efficient for capturing the spatial semantic information of the environment. Overall, each Gaussian includes 3D center position $\mu$, radius $r$, color $c=(r,g,b)$, opacity $\alpha$, 16-channel semantic feature embedding $e$, and is defined as standard Gaussian equation multiplied by opacity $\alpha$:
$g(x)=\alpha\exp\left(-\frac{\|x-\mu\|^{2}}{2r^{2}}\right).$ (1)
3D Gaussian Rendering.   Following Gaussian splatting [8], we project 3D Gaussians to 2D splats for high-fidelity differentiable rendering, allowing for explicit gradient flow for Gaussian scene optimization and pose estimation. Specifically, for splatting of semantic feature embedding, we first sort all Gaussians from front-to-back and then blend $N$ ordered points projecting to pixel $p=(u,v)$, obtaining 2D semantic feature maps $E(p)$:
$E(p)=\sum_{i\in N}e_{i}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)),$ (2)
where $g_{i}(p)$ is computed as shown in Eq.(1) with $\mu$ and $r$ representing the rendered 2D Gaussians in pixel plane:
$\mu_{\text{pix}}=K_{c}\frac{T_{k}\mu}{d},\ r_{\text{pix}}=\frac{fr}{d},\ \text{where}~{}d=(T_{k}\mu)_{z}.$ (3)
$K_{c}$ is calibrated camera intrinsic, $T_{k}$ is estimated camera pose at frame $k$, $f$ is known focal length, $d$ is the depth of the $i$-th Gaussian in camera coordinates.
For RGB and depth rendering, we follow the similar approach in Eq.(2):
$C(p)=\sum_{i\in N}c_{i}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)),\ \ D(p)=\sum_{i\in N}d_{i}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)),$ (4)
where $C(p)$ and $D(p)$ represents splatted 2D color and depth images respectively. Moreover, given camera pose, visibility information of Gaussian is required for mapping and tracking process, such as adding new Gaussian and loss construction. Therefore, following  [7], we render a silhouette image to determine visibility:
$Sil(p)=\sum_{i\in N}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)).$ (5)
Tracking Process.   During tracking, we keep 3D Gaussian parameters fixed and only optimize camera pose $T$ of current frame. Since the proximity between adjacent frames is relatively small, we assume a constant velocity model to obtain an initial pose estimation for new frame. The camera pose is then iteratively refined by optimization through loss construction between differentiable rendering from 3D Gaussian and camera observation within the visible silhouette.
Mapping Process.   Our system performs RGB mapping and semantic mapping simultaneously. Mapping process begins with the initialization of the scene representation, which is achieved by inverse transforming all pixels of the first frame to 3D coordinates and obtaining the initial 3D Gaussian representation. Then, when the overlap of coming frame with existing map rendering is less than half, we add a new Gaussian for incremental mapping. Figure 3(a) demonstrates 3D Gaussian semantic mapping process.
3.2 Loss Functions
For optimization of semantic scene representation, we utilize cross-entropy loss for constructing semantic loss $\mathcal{L}_{s}$. Furthermore, instead of using semantic loss solely, we introduce a feature-level loss $\mathcal{L}_{f}$ for higher-level semantic optimization guidance. Feature-level loss is obtained by constructing L1 loss between extracted features generated by Dinov2 [17]-based feature extractor and splatted features obtained from 3D Gaussian representation. Compared with semantic loss, feature loss provides direct guidance on intermediate features, leading to more robust and accurate semantic scene understanding.
We employ RGB loss $\mathcal{L}_{c}$ and depth loss $\mathcal{L}_{d}$ for optimization of scene color and geometry representation. $\mathcal{L}_{c}$ and $\mathcal{L}_{d}$ are both L1 loss constructed by comparing the RGB and depth splats with the input RGB-D frame. These loss functions are then utilized for mapping and pose estimation.
In mapping process, we construct loss over all rendered pixels for 3D Gaussian scene optimization. Moreover, we add SSIM term to RGB loss following [8]. The complete loss function for mapping is a weighted sum of the above losses:
$\mathcal{L}_{\text{mapping}}=\sum_{p\in P_{M}}{(\lambda_{f_{m}}\mathcal{L}_{f}(p)+\lambda_{s_{m}}\mathcal{L}_{s}(p)+\lambda_{c_{m}}\mathcal{L}_{c}(p)+\lambda_{d_{m}}\mathcal{L}_{d}(p))},$ (6)
where $P_{M}$ represents the collection of all pixels in rendered image. $\lambda_{f_{m}}$, $\lambda_{s_{m}}$, $\lambda_{c_{m}}$, $\lambda_{d_{m}}$ are weighting coefficients.
During tracking, utilizing an overly constrained loss function for pose estimation can lead to less precise camera pose and increased processing time, which is unsuitable for real-time tracking. Therefore, loss function for tracking is constructed based only on a weighted sum of RGB loss and depth loss:
$\mathcal{L}_{\text{tracking}}=\sum_{p\in P_{T}}{(\lambda_{c_{t}}\mathcal{L}_{c}(p)+\lambda_{d_{t}}\mathcal{L}_{d}(p))},$ (7)
where $\lambda_{c_{t}}$, $\lambda_{d_{t}}$ are weighting coefficients in tracking process. $P_{T}$ represents pixels that are rendered from well-optimized part of 3D Gaussian map, which is area that rendered visibility silhouette $Sil(p)$ is greater than 0.99.
3.3 Semantic-informed Bundle Adjustment
Currently, existing radiance field-based semantic SLAM systems utilize the latest input RGB-D frame to construct RGB and depth loss for pose estimation, while optimization of scene representation is then performed using estimated camera pose and the latest frame. However, relying solely on single-frame constraint for pose optimization can lead to cumulative drift in the tracking process due to the lack of global consistency. Furthermore, using only single-frame information for scene representation optimization can result in updates to the scene that are inconsistent on a global semantic level. To address this problem, we propose semantic-informed bundle adjustment, as illustrated in Fig. 3(b), to achieve joint optimization of 3D Gaussian representation and camera poses by leveraging multi-view constraints and semantic associations.
In semantic-informed bundle adjustment, we leverage the consistency of multi-view semantics to establish constraints. Specifically, rendered semantic feature is warped to its co-visible keyframe $j$ using estimated relative pose $T_{i}^{j}$, and constructs loss with extracted semantic feature $F_{j}^{e}$ of keyframe $j$ to obtain $\mathcal{L}_{\text{BA-sem}}$:
$\mathcal{L}_{\text{BA-sem}}=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},e)-F_{j}^{e}|)},$ (8)
where $\mathcal{G}(T_{i},e)$ represents splatted semantic embedding from 3D Gaussian $\mathcal{G}$ using camera pose $T_{i}$. Moreover, to achieve geometry and appearance consistency, we also warp rendered RGB and depth to co-visible keyframes to construct loss with corresponding keyframes following the similar approach in Eq.(8):
$\mathcal{L}_{\text{BA-rgb}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},c)-F_{j}^{c}|)},$
$\mathcal{L}_{\text{BA-depth}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},d)-F_{j}^{d}|)},$ (9)
where $F_{j}^{c}$ and $F_{j}^{d}$ are color and depth of keyframe $j$ respectively. $\mathcal{G}(T_{i},c)$ and $\mathcal{G}(T_{i},d)$ represents rendered RGB and depth. Therefore, overall loss function $\mathcal{L}_{\text{BA}}$ for joint optimization of corresponding keyframe poses and 3D Gaussian scene representation is the weighted sum of the above losses:
$\mathcal{L}_{\text{BA}}=\lambda_{e}\mathcal{L}_{\text{BA-sem}}+\lambda_{c}\mathcal{L}_{\text{BA-rgb}}+\lambda_{d}\mathcal{L}_{\text{BA-depth}},$ (10)
where $\lambda_{e}$, $\lambda_{c}$, $\lambda_{d}$ are weighting coefficients. This design leverages fast rendering capability of 3D Gaussian to achieve joint optimization of scene representation and camera poses in real time. Furthermore, semantic-informed BA integrates consistency and correlation of multi-perspective semantic, geometry, and appearance information, resulting in low-drift tracking and consistent mapping.
Table 1: Quantitative comparison of tracking accuracy for our proposed SemGauss-SLAM with other dense visual SLAM and semantic SLAM methods on Replica dataset [21]. We utilize RMSE (cm) metric. To ensure more objectivity in the results, each scene is tested and averaged with three independent runs. Our work outperforms previous radiance field-based semantic SLAM methods across all scenes. Best results are highlighted as first, second.
Methods room0 room1 room2 office0 office1 office2 office3 office4 Avg.
Visual SLAM iMAP[23] 6.33 3.46 2.65 3.31 1.42 7.17 6.32 2.55 4.15
NICE-SLAM [33] 1.86 2.37 2.26 1.50 1.01 1.85 5.67 3.53 2.51
Vox-Fusion [29] 1.37 1.90 1.47 1.35 1.76 1.18 1.11 1.64 1.47
Co-SLAM [25] 0.72 0.85 1.02 0.69 0.56 2.12 1.62 0.87 1.06
ESLAM [6] 0.76 0.71 0.56 0.53 0.49 0.58 0.74 0.64 0.62
Point-SLAM [19] 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72 0.52
SplaTAM [7] 0.31 0.40 0.29 0.47 0.27 0.29 0.32 0.55 0.36
Semantic SLAM SNI-SLAM [32] 0.50 0.55 0.45 0.35 0.41 0.33 0.62 0.50 0.46
DNS SLAM [9] 0.49 0.46 0.38 0.34 0.35 0.39 0.62 0.60 0.45
SemGauss-SLAM (Ours) 0.26 0.42 0.27 0.34 0.17 0.32 0.36 0.49 0.33
4 Experiments
4.1 Experimental Setup
Datasets.   We evaluate the performance of SemGauss-SLAM on two datasets with semantic ground truth annotations, including 8 scenes on simulated dataset Replica [21] and 5 scenes on real-world dataset ScanNet [2].
Metrics.   To evaluate the SLAM system and rendering quality, we follow metrics from [19]. For mapping evaluation, we use Depth L1 (cm). For tracking accuracy evaluation, we utilize ATE RMSE (cm) [22]. Moreover, we use PSNR (dB), SSIM, and LPIPS for evaluating RGB image rendering performance. Semantic segmentation is evaluated with respect to mIoU accuracy.
Baselines.   We compare our method with the existing state of-the-art dense visual SLAM, including NeRF-based SLAM [23, 33, 25, 6, 19, 29] and 3D Gaussian SLAM [7]. For comparison of dense semantic SLAM performance, we consider NeRF-based semantic SLAM including SNI-SLAM [32], DNS SLAM [9], and NIDS-SLAM [4] as baseline.
Implementation Details.   We run SemGauss-SLAM on NVIDIA RTX 4090 GPU. For experimental settings, we perform mapping every 8 frames. The weighting coefficients of each loss are $\lambda_{f_{m}}=0.01$, $\lambda_{s_{m}}=0.01$, $\lambda_{c_{m}}=0.5$, $\lambda_{d_{m}}=1$ in mapping, $\lambda_{c_{t}}=0.5$ and $\lambda_{d_{t}}=1$ in tracking. Moreover, we set $\lambda_{e}=0.004$, $\lambda_{c}=0.5$, $\lambda_{c}=1$ in semantic-informed bundle adjustment. Please refer to the supplementary for further details of our implementation.
4.2 Experimental Results
Tracking Results.   As shown in Tab. 1 and 2, our method achieves up to 47% relative increase in tracking accuracy compared with other dense semantic SLAM methods. This improvement is attributed to semantic-informed bundle adjustment, which provides multi-view constraint leveraging consistency of semantic information to reduce accumulated drift in tracking process.
Table 2: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on ScanNet dataset [2] for tracking metric RMSE (cm). The results are an average of three independent runs. Best results are highlighted as first, second, third.
Methods scene0000 scene0059 scene0169 scene0181 scene0207 Avg.
NICE-SLAM [33] 12.00 14.00 10.90 13.40 6.20 11.30
Vox-Fusion [29] 68.84 24.18 27.28 23.30 9.41 30.60
Point-SLAM [19] 10.24 7.81 22.16 14.77 9.54 12.90
SplaTAM [7] 12.83 10.10 12.08 11.10 7.46 10.71
SemGauss-SLAM (Ours) 12.56 7.97 9.05 9.78 8.97 9.67
Rendering Quality Results.   Tab. 3 shows rendering quality on the input views of Replica dataset [21]. Our method achieves the best performance in PSNR, SSIM, and LPIPS compared with other dense semantic SLAM.
Table 3: Quantitative comparison of training view rendering performance on Replica dataset [21]. Our work outperforms other radiance field-based semantic SLAM methods on all three metrics across all scenes.
Methods Metrics room0 room1 room2 office0 office1 office2 office3 office4 Avg.
Visual SLAM NICE-SLAM [33] PSNR $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
SSIM $\uparrow$ 0.689 0.757 0.814 0.874 0.886 0.797 0.801 0.856 0.809
LPIPS $\downarrow$ 0.330 0.271 0.208 0.229 0.181 0.235 0.209 0.198 0.233
Vox-Fusion [29] PSNR $\uparrow$ 22.39 22.36 23.92 27.79 29.83 20.33 23.47 25.21 24.41
SSIM $\uparrow$ 0.683 0.751 0.798 0.857 0.876 0.794 0.803 0.847 0.801
LPIPS $\downarrow$ 0.303 0.269 0.234 0.241 0.184 0.243 0.213 0.199 0.236
Co-SLAM [25] PSNR $\uparrow$ 27.27 28.45 29.06 34.14 34.87 28.43 28.76 30.91 30.24
SSIM $\uparrow$ 0.910 0.909 0.932 0.961 0.969 0.938 0.941 0.955 0.939
LPIPS $\downarrow$ 0.324 0.294 0.266 0.209 0.196 0.258 0.229 0.236 0.252
ESLAM [6] PSNR $\uparrow$ 25.32 27.77 29.08 33.71 30.20 28.09 28.77 29.71 29.08
SSIM $\uparrow$ 0.875 0.902 0.932 0.960 0.923 0.943 0.948 0.945 0.929
LPIPS $\downarrow$ 0.313 0.298 0.248 0.184 0.228 0.241 0.196 0.204 0.239
SplaTAM [7] PSNR $\uparrow$ 32.86 33.89 35.25 38.26 39.17 31.97 29.70 31.81 34.11
SSIM $\uparrow$ 0.978 0.969 0.979 0.977 0.978 0.968 0.949 0.949 0.968
LPIPS $\downarrow$ 0.072 0.103 0.081 0.092 0.093 0.102 0.121 0.152 0.102
Semantic SLAM SNI-SLAM [32] PSNR $\uparrow$ 25.91 28.17 29.15 33.86 30.34 29.10 29.02 29.87 29.43
SSIM $\uparrow$ 0.885 0.910 0.938 0.965 0.927 0.950 0.950 0.952 0.935
LPIPS $\downarrow$ 0.307 0.292 0.245 0.182 0.225 0.238 0.192 0.198 0.235
SemGauss-SLAM (Ours) PSNR $\uparrow$ 32.55 33.92 35.15 39.18 39.87 32.97 31.60 35.00 35.03
SSIM $\uparrow$ 0.979 0.979 0.987 0.989 0.990 0.979 0.972 0.978 0.982
LPIPS $\downarrow$ 0.055 0.054 0.045 0.048 0.050 0.069 0.078 0.093 0.062
Reconstruction Results.   As shown in Tab. 4, our method achieves up to 77.3% increase in reconstruction accuracy compared with other radiance field-based methods. This enhancement in performance is due to the incorporation of semantic-informed BA process, achieving joint optimization for both camera poses and scene representation. Specifically, our method exploits the geometric consistency of co-visible frames to construct geometry constraints for more precise reconstruction.
Table 4: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on Replica [21] for reconstruction metric Depth L1 (cm). The results are an average of three independent runs. Our method outperforms other methods by up to 77.3%.
Methods room0 room1 room2 office0 office1 office2 office3 office4 Avg.
Visual SLAM NICE-SLAM [33] 1.81 1.44 2.04 1.39 1.76 8.33 4.99 2.01 2.97
Vox-Fusion [29] 1.09 1.90 2.21 2.32 3.40 4.19 2.96 1.61 2.46
Co-SLAM [25] 1.05 0.85 2.37 1.24 1.48 1.86 1.66 1.54 1.51
ESLAM [6] 0.73 0.74 1.26 0.71 1.02 0.93 1.03 1.18 0.95
Semantic SLAM SNI-SLAM [32] 0.55 0.58 0.87 0.55 0.97 0.89 0.75 0.97 0.77
SemGauss-SLAM (Ours) 0.54 0.46 0.43 0.29 0.22 0.51 0.98 0.56 0.50
Semantic Segmentation Results.   As shown in Tab. 5, our work outperforms other dense semantic SLAM methods by up to 26% on mIoU metric and achieves 95% mIoU accuracy. Such enhancement attributes to the integration of semantic feature embedding into 3D Gaussian for enriched semantic representation, and semantic feature-level loss for direct guidance of semantic optimization. Moreover, our proposed semantic-informed BA also contributes to high semantic precision, as it leverages multiple co-visible frames to construct a globally consistent semantic map, thereby achieving high-precision semantic representation.
Table 5: Quantitative comparison of input views semantic segmentation performance on Replica [21] for semantic metric mIoU(%). For one scene, we calculate mIoU between splatted semantic maps and ground truth labels, which is calculated every 4 frames, to obtain average mIoU. Our method achieves highest semantic accuracy across all scenes.
Methods room0 room1 room2 office0 office1 office2 office3 office4
NIDS-SLAM [4] 82.45 84.08 76.99 85.94
DNS SLAM [9] 88.32 84.90 81.20 84.66
SNI-SLAM [32] 88.42 87.43 86.16 87.63 78.63 86.49 74.01 80.22
SemGauss-SLAM (Ours) 92.81 94.10 94.72 95.23 90.11 94.93 92.93 94.82
Table 6: Quantitative comparison of semantic novel view synthesis performance on Replica [21] for semantic metric mIoU(%). For one scene, we randomly choose 100 new viewpoints for semantic novel view synthesis evaluation. Our method outperforms SNI-SLAM [32] across all scenes, showing excellent 3D semantic mapping accuracy.
Methods room0 room1 room2 office0 office1 office2 office3 office4 Avg.
SNI-SLAM [32] 51.20 50.10 54.80 70.21 63.47 58.91 64.31 71.04 60.51
SemGauss-SLAM (Ours) 89.63 84.72 86.51 93.60 89.57 92.80 92.41 92.10 90.17
Novel View Semantic Evaluation Results.   As shown in Tab. 6, our method achieves up to 75% increase in semantic segmentation accuracy for semantic novel view synthesis compared with existing dense semantic SLAM. By introducing 3D Gaussian feature embedding for semantic representation, our method enables continuous semantic modeling. This modeling is crucial for generating semantically coherent scenes, as it reduces the occurrence of sharp transitions that can cause inconsistencies, ensuring semantic representation accuracy from different viewpoints for high-precision novel view semantic synthesis.
Figure 4: Qualitative comparison on rendering quality of our method and baseline. We visualize 4 selected scenes of Replica dataset[21] and details are highlighted with red color boxes. Our method achieves photo-realistic rendering quality compared with other methods, especially in areas with rich textural information.
Visualization.   Fig. 4 shows rendering quality comparison of 4 scenes with interesting regions highlighted with colored boxes. For small objects, such as vases and items on the table, other methods fail to reconstruct them clearly. Our method leverages the high-quality rendering capability of 3D Gaussian representation and introduces semantic-informed BA to achieve detailed geometric reconstruction results. Specifically, our method enhances multi-view geometry consistency by BA to ensure that the reconstructed geometry aligns accurately across all observed viewpoints, leading to precise geometry reconstruction. As shown in Fig. 5, our method achieves superior novel view semantic segmentation accuracy compared with the baseline SNI-SLAM [32]. It can be observed that SNI-SLAM struggles with segmenting ceilings in novel view synthesis as they are less frequently observed during the mapping process. Consequently, the ceiling features are poorly modeled in the semantic scene representation. Our method introduces semantic-informed BA to construct multi-view constraints, enabling areas that are sparsely observed to effectively utilize the limited information from co-visible frames, thus establishing sufficient constraints for accurate semantic reconstruction.
Figure 5: Qualitative comparison on semantic novel view synthesis of our method and baseline semantic SLAM method SNI-SLAM [32] on 4 scenes of Replica[21].
Figure 6: Semantic rendering results and ground truth labels of feature-level loss ablation on two scenes of Replica [21].
4.3 Ablation Study
We perform ablation study on three scenes of Replica dataset[21] in Tab. 7 to validate the effectiveness of feature-level loss and semantic-informed BA in SemGauss-SLAM. Moreover, ablation of semantic-informed BA component is conducted in Tab. 8.
Feature-level Loss.   Tab. 7 shows that incorporating feature-level loss significantly enhances semantic segmentation performance, while having little effect on tracking and geometric reconstruction. This result occurs because feature-level loss influences only the optimization of semantic features, without affecting the optimization of geometry and pose estimation. As shown in Fig. 6, utilizing feature-level loss can achieve improved boundary segmentation and finer segmentation of small objects. This enhancement occurs because feature loss compels the scene representation to capture high-dimensional and direct information within the feature space, leading to the capability of distinguishing intricate details and subtle variations within the scene.
Semantic-informed BA.   Tab. 7 shows that introducing semantic-informed BA leads to enhancements in tracking, reconstruction, and semantic segmentation. This improvement is due to the joint optimization of camera poses and scene representation, which is informed by multi-view constraints. As shown in Tab. 8, lacking semantic constraint leads to a significant decrease in tracking performance compared with the absence of color and depth constraints. Such result suggests that multi-view semantic constraints provide more comprehensive and accurate information, due to the consistency of semantics across multiple perspectives. Moreover, the absence of semantic constraint can lead to reduced semantic precision, while lacking color and depth constraints results in decreased reconstruction accuracy.
Table 7: Ablation study of our contributions on three scenes of Replica [21].
Methods room0 room1 office1
RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
w/o feature-level loss 0.26 0.54 83.60 0.42 0.46 86.10 0.17 0.22 80.13
w/o semantic-informed BA 0.35 0.70 90.10 0.51 0.60 91.08 0.21 0.30 87.73
SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.42 0.46 94.10 0.17 0.22 90.11
Table 8: Ablation study of semantic-informed BA on three scenes of Replica [21]. (w/o semantic) without adding semantic constraint loss $\mathcal{L}_{\text{BA-sem}}$; (w/o RGB and depth) without adding RGB and depth constraint loss $\mathcal{L}_{\text{BA-rgb}}$ and $\mathcal{L}_{\text{BA-depth}}$.
Methods room0 room1 office1
RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
w/o semantic 0.35 0.66 90.15 0.50 0.49 91.50 0.20 0.24 87.95
w/o RGB and depth 0.31 0.70 92.01 0.48 0.59 93.90 0.18 0.28 89.60
SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.42 0.46 94.10 0.17 0.22 90.11
5 Conclusion
We propose SemGauss-SLAM, a novel dense semantic SLAM system utilizing 3D Gaussian representation that enables dense visual mapping, robust camera tracking, and 3D semantic mapping of the whole scene. We incorporate semantic feature embedding into 3D Gaussian for Gaussian semantic representation, leading to real-time dense semantic mapping. Moreover, we propose feature-level loss for 3D Gaussian scene optimization to achieve accurate semantic representation. In addition, we introduce semantic-informed BA that enables joint optimization of camera poses and 3D Gaussian representation by establishing multi-view semantic constraints, resulting in low-drift tracking and precise mapping.

### v3

@@
 Abstract
-We propose SemGauss-SLAM, the first semantic SLAM system utilizing 3D Gaussian representation, that enables accurate 3D semantic mapping, robust camera tracking, and high-quality rendering in real-time.
+We propose SemGauss-SLAM, a dense semantic SLAM system utilizing 3D Gaussian representation, that enables accurate 3D semantic mapping, robust camera tracking, and high-quality rendering simultaneously.
 In this system, we incorporate semantic feature embedding into 3D Gaussian representation, which effectively encodes semantic information within the spatial layout of the environment for precise semantic scene representation.
 Furthermore, we propose feature-level loss for updating 3D Gaussian representation, enabling higher-level guidance for 3D Gaussian optimization.
-In addition, to reduce cumulative drift and improve reconstruction accuracy, we introduce semantic-informed bundle adjustment leveraging semantic associations for joint optimization of 3D Gaussian representation and camera poses, leading to more robust tracking and consistent mapping.
-Our SemGauss-SLAM method demonstrates superior performance over existing dense semantic SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in novel-view semantic synthesis and 3D semantic mapping.
-The source code will be released upon acceptance.
-Keywords:
-Dense Semantic SLAM Semantic-informed Bundle Adjustment 3D Gaussian Semantic Representation 3D Reconstruction
+In addition, to reduce cumulative drift in tracking and improve semantic reconstruction accuracy, we introduce semantic-informed bundle adjustment leveraging multi-frame semantic associations for joint optimization of 3D Gaussian representation and camera poses, leading to low-drift tracking and accurate mapping.
+Our SemGauss-SLAM method demonstrates superior performance over existing radiance field-based SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in high-precision semantic segmentation and dense semantic mapping.
 1 Introduction
-Dense semantic Simultaneous Localization and Mapping (SLAM) is a fundamental challenge for robotic systems [18, 15] and autonomous driving [1, 10].
+Dense semantic Simultaneous Localization and Mapping (SLAM) is a fundamental challenge for robotic systems rosinol2020kimera ; mccormac2017semanticfusion and autonomous driving bao2022semantic ; lianos2018vso ; wang2022sfgan ; jiang20223d ; jiang20243dsflabelling ; jiang2022ffpa ; feng2024rogs .
 It integrates semantic understanding of the environment into dense map reconstruction and performs pose estimation simultaneously.
-Traditional semantic SLAM has limitations including its inability to predict unknown areas and the demand for considerable map storage [11].
-Subsequent semantic SLAM based on Neural Radiance Fields (NeRF) [16] methods [32, 9] have addressed these drawbacks, but suffers from inefficient per-pixel raycasting rendering.
-Recently, a novel radiance field based on 3D Gaussian [8] has demonstrated remarkable capability in scene representation, enabling high-quality and efficient rendering via splatting.
-Following the advantages of 3D Gaussian representation, 3D Gaussian SLAM methods [27, 7, 31, 13] have been developed to achieve photo-realistic mapping.
-However, existing 3D Gaussian SLAM systems focus on visual mapping to obtain RGB maps, where color information alone is insufficient for downstream tasks such as navigation.
-Meanwhile, current semantic SLAM methods based on NeRF are prone to cumulative drift, leading to degraded SLAM accuracy.
-Therefore, developing a low-drift semantic SLAM system based on radiance field is essential and challenging.
+Traditional dense semantic SLAM has limitations including its inability to predict unknown areas liu2023unsupervised .
+Subsequent semantic SLAM based on Neural Radiance Fields (NeRF) mildenhall2021nerf methods zhu2023sni ; li2023dns ; haghighi2023neural have addressed these drawbacks, but suffer from inefficient per-pixel raycasting rendering for real-time optimization of implicit scene representation and low-quality novel view semantic representation.
+Recently, a novel radiance field based on 3D Gaussian Splatting (3DGS) kerbl20233d has demonstrated remarkable capability in scene representation, enabling high-quality and efficient rendering via splatting.
+Following the advantages of 3D Gaussian representation, 3DGS-based SLAM methods yan2023gs ; keetha2023splatam ; yugay2023gaussian ; matsuki2023gaussian ; huang2023photo have been developed to achieve photo-realistic mapping.
+However, most existing 3DGS SLAM systems focus on visual mapping to obtain RGB maps, where color information alone is insufficient for downstream tasks such as navigation.
+Moreover, current semantic NeRF-based SLAM methods suffer from cumulative drift in tracking, leading to degraded SLAM accuracy.
+Additionally, these methods struggle to achieve accurate semantic segmentation from novel viewpoints, indicating limited precision in online-reconstructed 3D semantic representation.
+Therefore, developing a low-drift and high-precision semantic SLAM system based on radiance field is essential and challenging.
 Figure 1: Our SemGauss-SLAM incorporates semantic feature embedding into 3D Gaussian representation to perform dense semantic SLAM.
@@
 Moreover, semantic mapping is visualized using semantic feature embedding, showing 3D semantic modeling of the scene.
-For radiance field-based semantic SLAM, there are two challenges: 1) Existing semantic SLAM methods struggle to achieve real-time 3D semantic mapping from 2D semantic information without predefined bounds, which is a fundamental requirement for semantic SLAM system.
-2) As the progression of tracking, pose estimation is prone to cumulative drift, leading to a decrease in SLAM accuracy.
-For the first challenge, SNI-SLAM [32] utilizes feature collaboration to achieve semantic mapping but requires specific scene bounds.
-Moreover, solely utilizing rendering results for semantic representation optimization without initial value is computationally inefficient.
-In this paper, we leverage the explicit structure of 3D Gaussian representation and propagate extracted 2D semantic features directly to 3D Gaussian as the initial value for achieving unbounded mapping and efficient 3D semantic scene optimization.
-This design enables faster transfer from 2D semantic information to 3D semantic map.
-In addition, we introduce 2D feature-level loss to guide scene representation optimization at a higher-level, thereby accelerating the convergence of 3D Gaussian optimization.
-For the second challenge, NICE-SLAM [33] performs local bundle adjustment (BA) for joint optimization of pose and scene representation to reduce accumulated drift.
-However, it only utilizes appearance and geometry constraints, lacking adequate semantic constraints for semantic SLAM optimization.
-To address this challenge, we leverage semantic associations among co-visible frames and propose semantic-informed bundle adjustment for joint optimization of camera poses and 3D Gaussian representation.
+To overcome this challenge, we introduce a novel dense semantic SLAM system based on 3D Gaussian Splatting with semantic-informed bundle adjustment, which enables high accuracy in tracking, dense mapping, and semantic segmentation from novel view.
+Different from concurrent work SGS-SLAM li2024sgs , which uses color for semantic representation to achieve 3DGS semantic SLAM, our method integrates semantic feature embedding into 3D Gaussian for semantic modeling, resulting in a more discriminative and comprehensive understanding of the scene.
+Moreover, by leveraging the explicit 3D structure of 3D Gaussian representation, our embedding-based semantic Gaussian representation is capable of capturing the spatial distribution of semantic information, achieving high-accuracy novel view semantic segmentation.
+Furthermore, to reduce accumulated drift in radiance field-based semantic SLAM, we leverage semantic associations among co-visible frames and propose semantic-informed bundle adjustment (BA) for joint optimization of camera poses and 3D Gaussian representation.
 This design exploits the consistency of multi-view semantics for establishing constraints, which enables the reduction of cumulative drift in tracking and enhanced semantic mapping precision.
 Overall, we provide the following contributions:
-  • We present SemGauss-SLAM, the first 3D Gaussian semantic SLAM system, which can achieve accurate semantic mapping and photo-realistic reconstruction.
-We incorporate semantic feature embedding into 3D Gaussian for precisely constructing semantic maps.
-Moreover, feature-level loss is introduced to provide higher-level guidance for 3D Gaussian optimization.
+  • We present SemGauss-SLAM, a dense semantic SLAM system based on 3D Gaussian, which can achieve accurate semantic mapping, photo-realistic reconstruction, and robust tracking.
+We incorporate semantic feature embedding into 3D Gaussian to achieve precise semantic scene representation.
+Moreover, feature-level loss is introduced to provide higher-level guidance for 3D Gaussian optimization, leading to high-precision scene optimization results.
   • We perform semantic-informed bundle adjustment by leveraging multi-view semantic constraints for joint optimization of camera poses and 3D Gaussian representation, achieving low-drift tracking and accurate semantic mapping.
-  • We conduct extensive evaluations on two challenging datasets, Replica [21] and ScanNet [2], to demonstrate our method achieves state-of-the-art performance compared with existing radiance field-based SLAM in mapping, tracking, semantic segmentation and novel-view synthesis.
+  • We conduct extensive evaluations on two challenging datasets, Replica straub2019replica and ScanNet dai2017scannet , to demonstrate our method achieves state-of-the-art performance compared with existing radiance field-based SLAM in mapping, tracking, semantic segmentation and novel-view synthesis.
 2 Related Work
-Semantic SLAM can be divided into two main categories based on the form of scene representation: traditional semantic SLAM and neural implicit semantic SLAM.
-Traditional semantic SLAM employs explicit 3D representation such as surfels [15], mesh [18, 24, 5] and Truncated Signed Distance Fields (TSDF) [20, 14, 3].
-Neural implicit semantic SLAM [4, 9, 32] utilizes implicit representation, such as feature grid [9] and feature plane [32].
 Traditional Semantic SLAM.
-SemanticFusion [15] uses surfel representation and employs Conditional Random Field (CRF) for updating class probability distribution incrementally.
-Fusion++ [14] performs object-level SLAM where each object is reconstructed within its own TSDF volume and segmented based on estimated foreground probability.
-Kimera [18] utilizes visual-inertial odometry for pose estimation and generates dense semantic mesh maps.
-However, these explicit scene modeling methods not only require high storage space but also fail to achieve high-fidelity and complete reconstruction.
+Traditional semantic SLAM employs explicit 3D representation such as surfels mccormac2017semanticfusion , mesh rosinol2020kimera ; tian2022kimera ; hughes2022hydra and Truncated Signed Distance Fields (TSDF) schmid2022panoptic ; mccormac2018fusion++ ; grinvald2019volumetric for dense semantic mapping.
+SemanticFusion mccormac2017semanticfusion uses surfel representation and employs Conditional Random Field (CRF) for updating class probability distribution incrementally.
+Fusion++ mccormac2018fusion++ performs object-level SLAM where each object is reconstructed within its own TSDF volume and segmented based on estimated foreground probability.
+Kimera rosinol2020kimera utilizes visual-inertial odometry for pose estimation and generates dense semantic mesh maps.
+However, these explicit scene modeling methods not only require high storage space but also fail to achieve high-fidelity and complete 3D reconstruction.
 Neural Implicit SLAM.
-iMAP [23] first achieves real-time mapping and tracking utilizing a single MLP network for scene representation.
-To overcome oversmoothed scene reconstruction and improve scalability, NICE-SLAM [33] adopts hierarchical feature grid representation.
-Following this, several works [29, 6, 25, 19] introduce more efficient scene representation, such as hash-based feature grid and feature plane, to achieve more accurate SLAM performance.
-For neural implicit semantic SLAM, DNS SLAM [9] utilizes 2D semantic priors and integrates multi-view geometry constraints for semantic reconstruction.
-SNI-SLAM [32] introduces feature collaboration and one-way correlation decoder for improved scene representation in semantic mapping.
+Neural implicit representation mildenhall2021nerf has shown promising capability in scene reconstruction for dense mapping.
+iMAP imap first achieves real-time mapping and tracking utilizing a single MLP network for scene representation.
+To overcome over-smoothed scene reconstruction and improve scalability, NICE-SLAM zhu2022nice adopts hierarchical feature grid representation.
+Following this, several works yang2022vox ; eslam ; wang2023co ; sandstrom2023point ; kong2023vmap ; hu2024cp ; deng2023plgslam introduce more efficient scene representation, such as hash-based feature grid and feature plane, to achieve more accurate SLAM performance.
+For neural implicit semantic SLAM, DNS SLAM li2023dns utilizes 2D semantic priors and integrates multi-view geometry constraints for semantic reconstruction.
+SNI-SLAM zhu2023sni introduces feature collaboration and one-way correlation decoder for improved scene representation in semantic mapping.
 However, these methods require specific scene bounds for mapping and suffer from cumulative drift in pose estimation.
-In this paper, we leverage the explicit structure of 3D Gaussian for unbounded mapping, and perform semantic-informed bundle adjustment utilizing multi-frame semantic constraints for conducting low-drift, high-quality dense semantic SLAM.
+In this paper, we leverage the explicit structure of 3D Gaussian for unbounded mapping and perform semantic-informed bundle adjustment utilizing multi-frame semantic constraints for conducting low-drift and high-quality dense semantic SLAM.
 3D Gaussian Splatting SLAM.
-3D Gaussian [8] emerges as a promising 3D scene representation using a set of 3D Gaussians with position, anisotropic covariance, opacity, and color.
-This representation is capable of quick differential rendering through splatting and has a wide range of applications in dynamic scene modeling [12, 26, 28] and scene editing [34, 30].
-Our main focus is on 3D Gaussian SLAM [27, 7, 13, 31].
+3D Gaussian Splatting kerbl20233d emerges as a promising 3D scene representation using a set of 3D Gaussians with learnable properties, including 3D center position, anisotropic covariance, opacity, and color.
+This representation is capable of quick differential rendering through splatting and has a wide range of applications in dynamic scene modeling luiten2023dynamic ; wu20234d ; yang2023deformable ; yan2024street ; zhou2023drivinggaussian , semantic segmentation cen2023segment ; ye2023gaussian ; zhou2023feature , scene editing zhuang2024tip ; fang2023gaussianeditor ; huang2023point .
+Our main focus is on 3D Gaussian SLAM yan2023gs ; keetha2023splatam ; matsuki2023gaussian ; yugay2023gaussian .
 These works emerge concurrently and all perform dense visual SLAM by leveraging scene geometry and appearance modeling capabilities of 3D Gaussian representation.
-SplaTAM [7] introduces silhouette-guided optimization to facilitate structured map expansion for dense mapping of visual SLAM.
-Gaussian Splatting SLAM [13] performs novel Gaussian insertion and pruning for monocular SLAM.
-However, the capability of 3D Gaussian representation extends well beyond appearance and geometry modeling of scene, as it can be augmented to perform semantic scene understanding.
-In this paper, we incorporate semantic feature embedding into 3D Gaussian for 3D semantic scene modeling to achieve high-precision dense semantic SLAM.
+SplaTAM keetha2023splatam introduces silhouette-guided optimization to facilitate structured map expansion for dense mapping of visual SLAM.
+Gaussian Splatting SLAM matsuki2023gaussian performs novel Gaussian insertion and pruning for monocular SLAM.
+However, the capability of 3D Gaussian representation extends well beyond appearance and geometry modeling of the scene, as it can be augmented to perform semantic scene understanding.
+The concurrent work SGS-SLAM li2024sgs introduces semantic color associated with the Gaussian for semantic representation.
+However, this color-based semantic modeling method ignores the higher-level information inherent in semantics, which is insufficient for semantic representation of the environment.
+To address this limitation, we incorporate semantic feature embedding into 3D Gaussian for 3D semantic scene modeling to achieve high-precision dense semantic SLAM.
 Figure 2: An overview of SemGauss-SLAM.
@@
 RGB images are fed into feature extractor to obtain semantic features.
-These features are then categorized by pretrained classifier to attain semantic labels.
+These features are then categorized by a pretrained classifier to attain semantic labels.
 Then, semantic features, semantic labels, along with the input RGB and depth data serve as supervision signals.
 In the meantime, semantic features and input RGB-D data propagate to 3D Gaussian blobs as initial properties of Gaussian representation.
-Rendered semantic feature, RGB, and depth are obtained from 3D Gaussian splatting, while rendered semantic label is attained by classifying rendered feature.
+Rendered semantic features, RGB, and depth are obtained from 3D Gaussian splatting, while rendered semantic labels are attained by classifying rendered features.
 Supervision and rendered information are utilized for loss construction to optimize camera poses and 3D Gaussian representation.
@@
 3 Method
-SemGauss-SLAM is the first RGB-D dense semantic SLAM based on 3D Gaussian splatting.
 The overview of SemGauss-SLAM is shown in Fig.
@@
 Sec.
-3.1 introduces semantic Gaussian representation for dense semantic mapping, as well as the process of 3D Gaussian-based semantic SLAM.
+3.1 introduces semantic Gaussian representation for dense semantic mapping, as well as the process of 3DGS-based semantic SLAM.
 Sec.
-3.2 introduces the loss functions in the mapping and tracking process.
+3.2 introduces loss functions in the mapping and tracking process.
 Sec.
-3.3 presents semantic-informed bundle adjustment to reduce accumulated drift in tracking and improve semantic mapping accuracy.
-(a) 3D Gaussian Semantic Mapping
-(b) Semantic-informed Bundle Adjustment
-Figure 3: (a) 3D Gaussian semantic mapping process that utilizes 3D Gaussian blobs with semantic feature embedding for visualization. (b) Illustration of semantic-informed bundle adjustment.
+3.3 presents semantic-informed bundle adjustment to reduce accumulated drift in tracking and improve mapping accuracy.
 3.1 3D Gaussian Semantic Mapping and Tracking
 Semantic Gaussian Representation.
-We utilize a set of Gaussians with specific properties for scene representation.
-To simplify scene representation for the SLAM process, we use isotropy Gaussians.
-Considering semantic representation, which has never been explored in other Gaussian SLAM, we introduce a new parameter, semantic feature embedding, to each Gaussian for semantic representation.
-To achieve real-time semantic mapping, it is crucial for 3D Gaussian representation, augmented by semantic feature embedding, to converge rapidly during the mapping optimization process.
-Therefore, instead of initializing these feature embeddings randomly, we propagate 2D semantic features extracted from images to 3D Gaussian as the initial values to achieve faster convergence of Gaussian semantic feature optimization.
-Moreover, integrating feature embedding into 3D Gaussian makes Gaussian semantic representation compact and efficient for capturing the spatial semantic information of the environment.
+We utilize a set of isotropy Gaussians with specific properties for scene representation in SLAM.
+To achieve semantic Gaussian mapping, we introduce a new parameter, semantic feature embedding, to each Gaussian for semantic representation.
+This design enables compact and efficient Gaussian semantic representation to capture the spatial semantic information of the environment.
+Moreover, it is crucial for 3D Gaussian representation, augmented by semantic feature embedding, to converge rapidly during optimization process for real-time mapping.
+Therefore, instead of initializing these feature embeddings randomly, we propagate 2D semantic features extracted from images to 3D Gaussian as the initial values to achieve faster convergence of semantic Gaussian optimization.
+In this work, we use a universal feature extractor Dinov2 oquab2023dinov2 , followed by a pretrained classifier to construct the segmentation network.
 Overall, each Gaussian includes 3D center position $\mu$, radius $r$, color $c=(r,g,b)$, opacity $\alpha$, 16-channel semantic feature embedding $e$, and is defined as standard Gaussian equation multiplied by opacity $\alpha$:
@@
 3D Gaussian Rendering.
-Following Gaussian splatting [8], we project 3D Gaussians to 2D splats for high-fidelity differentiable rendering, allowing for explicit gradient flow for Gaussian scene optimization and pose estimation.
-Specifically, for splatting of semantic feature embedding, we first sort all Gaussians from front-to-back and then blend $N$ ordered points projecting to pixel $p=(u,v)$, obtaining 2D semantic feature maps $E(p)$:
+Following Gaussian splatting kerbl20233d , we project 3D Gaussians to 2D splats for high-fidelity differentiable rendering, which facilitates explicit gradient flow for Gaussian scene optimization and pose estimation.
+Specifically, for splatting of semantic feature embedding, we first sort all Gaussians in order of depth from near to far and then blend $N$ ordered points projecting to pixel $p=(u,v)$, obtaining 2D semantic feature maps $E(p)$:
 $E(p)=\sum_{i\in N}e_{i}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)),$ (2)
@@
 where $C(p)$ and $D(p)$ represents splatted 2D color and depth images respectively.
-Moreover, given camera pose, visibility information of Gaussian is required for mapping and tracking process, such as adding new Gaussian and loss construction.
-Therefore, following  [7], we render a silhouette image to determine visibility:
+Moreover, given the camera pose, visibility information of Gaussian is required for mapping and tracking process, such as adding new Gaussian and loss construction.
+Therefore, following  keetha2023splatam , we render a silhouette image to determine visibility:
 $Sil(p)=\sum_{i\in N}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)).$ (5)
@@
 During tracking, we keep 3D Gaussian parameters fixed and only optimize camera pose $T$ of current frame.
-Since the proximity between adjacent frames is relatively small, we assume a constant velocity model to obtain an initial pose estimation for new frame.
+Since the proximity between adjacent frames is relatively small, we assume a constant velocity model to obtain an initial pose estimation for the new frame.
 The camera pose is then iteratively refined by optimization through loss construction between differentiable rendering from 3D Gaussian and camera observation within the visible silhouette.
@@
 Mapping process begins with the initialization of the scene representation, which is achieved by inverse transforming all pixels of the first frame to 3D coordinates and obtaining the initial 3D Gaussian representation.
-Then, when the overlap of coming frame with existing map rendering is less than half, we add a new Gaussian for incremental mapping.
-Figure 3(a) demonstrates 3D Gaussian semantic mapping process.
+Then, when the overlap of coming frame with rendering of the existing map is less than half, we add a new Gaussian for incremental mapping.
 3.2 Loss Functions
@@
 Furthermore, instead of using semantic loss solely, we introduce a feature-level loss $\mathcal{L}_{f}$ for higher-level semantic optimization guidance.
-Feature-level loss is obtained by constructing L1 loss between extracted features generated by Dinov2 [17]-based feature extractor and splatted features obtained from 3D Gaussian representation.
+Feature-level loss is obtained by constructing L1 loss between extracted features generated by Dinov2 oquab2023dinov2 -based feature extractor and splatted features obtained from 3D Gaussian representation.
 Compared with semantic loss, feature loss provides direct guidance on intermediate features, leading to more robust and accurate semantic scene understanding.
@@
 In mapping process, we construct loss over all rendered pixels for 3D Gaussian scene optimization.
-Moreover, we add SSIM term to RGB loss following [8].
+Moreover, we add SSIM term to RGB loss following kerbl20233d .
 The complete loss function for mapping is a weighted sum of the above losses:
-$\mathcal{L}_{\text{mapping}}=\sum_{p\in P_{M}}{(\lambda_{f_{m}}\mathcal{L}_{f}(p)+\lambda_{s_{m}}\mathcal{L}_{s}(p)+\lambda_{c_{m}}\mathcal{L}_{c}(p)+\lambda_{d_{m}}\mathcal{L}_{d}(p))},$ (6)
-where $P_{M}$ represents the collection of all pixels in rendered image. $\lambda_{f_{m}}$, $\lambda_{s_{m}}$, $\lambda_{c_{m}}$, $\lambda_{d_{m}}$ are weighting coefficients.
-During tracking, utilizing an overly constrained loss function for pose estimation can lead to less precise camera pose and increased processing time, which is unsuitable for real-time tracking.
+$\mathcal{L}_{\text{mapping}}=\sum_{p\in P_{M}}{(\lambda_{f_{m}}\mathcal{L}_{f}(p)+\lambda_{s_{m}}\mathcal{L}_{s}(p)+\lambda_{c_{m}}\mathcal{L}_{c}(p)+\lambda_{d_{m}}\mathcal{L}_{d}(p))}.$ (6)
+$P_{M}$ represents a set of all pixels in rendered image. $\lambda_{f_{m}}$, $\lambda_{s_{m}}$, $\lambda_{c_{m}}$, $\lambda_{d_{m}}$ are weighting coefficients.
+During tracking, utilizing an overly constrained loss function for pose estimation can lead to decreased accuracy in camera pose and increased processing time.
 Therefore, loss function for tracking is constructed based only on a weighted sum of RGB loss and depth loss:
@@
 3.3 Semantic-informed Bundle Adjustment
-Currently, existing radiance field-based semantic SLAM systems utilize the latest input RGB-D frame to construct RGB and depth loss for pose estimation, while optimization of scene representation is then performed using estimated camera pose and the latest frame.
-However, relying solely on single-frame constraint for pose optimization can lead to cumulative drift in the tracking process due to the lack of global consistency.
-Furthermore, using only single-frame information for scene representation optimization can result in updates to the scene that are inconsistent on a global semantic level.
-To address this problem, we propose semantic-informed bundle adjustment, as illustrated in Fig.
-3(b), to achieve joint optimization of 3D Gaussian representation and camera poses by leveraging multi-view constraints and semantic associations.
-In semantic-informed bundle adjustment, we leverage the consistency of multi-view semantics to establish constraints.
-Specifically, rendered semantic feature is warped to its co-visible keyframe $j$ using estimated relative pose $T_{i}^{j}$, and constructs loss with extracted semantic feature $F_{j}^{e}$ of keyframe $j$ to obtain $\mathcal{L}_{\text{BA-sem}}$:
-$\mathcal{L}_{\text{BA-sem}}=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},e)-F_{j}^{e}|)},$ (8)
+Currently, existing radiance field-based semantic SLAM systems utilize the latest input RGB-D frame to construct RGB and depth loss for pose estimation.
+Subsequently, the scene representation of these SLAM systems is optimized using the estimated pose and the latest frame.
+However, relying solely on single-frame constraint for pose optimization can lead to cumulative drift in the tracking process due to the absence of global constraints.
+Furthermore, using only single-frame information to optimize scene representation can result in globally inconsistent updates to the scene on the semantic level.
+To address this problem, we propose semantic-informed bundle adjustment to achieve joint optimization of 3D Gaussian representation and camera poses by leveraging multi-view constraints and semantic associations.
+In semantic-informed BA, we leverage the consistency of multi-view semantics to establish constraints.
+Specifically, rendered semantic feature is warped to its co-visible frame $j$ using estimated relative pose transformation $T_{i}^{j}$, and constructs loss with rendered semantic feature $\mathcal{G}(T_{j},e)$ of frame $j$ to obtain $\mathcal{L}_{\text{BA-sem}}$:
+$\mathcal{L}_{\text{BA-sem}}=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},e)-\mathcal{G}(T_{j},e)|)},$ (8)
 where $\mathcal{G}(T_{i},e)$ represents splatted semantic embedding from 3D Gaussian $\mathcal{G}$ using camera pose $T_{i}$.
-Moreover, to achieve geometry and appearance consistency, we also warp rendered RGB and depth to co-visible keyframes to construct loss with corresponding keyframes following the similar approach in Eq.(8):
-$\mathcal{L}_{\text{BA-rgb}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},c)-F_{j}^{c}|)},$
-$\mathcal{L}_{\text{BA-depth}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},d)-F_{j}^{d}|)},$ (9)
-where $F_{j}^{c}$ and $F_{j}^{d}$ are color and depth of keyframe $j$ respectively.
-$\mathcal{G}(T_{i},c)$ and $\mathcal{G}(T_{i},d)$ represents rendered RGB and depth.
-Therefore, overall loss function $\mathcal{L}_{\text{BA}}$ for joint optimization of corresponding keyframe poses and 3D Gaussian scene representation is the weighted sum of the above losses:
+Moreover, to achieve geometry and appearance consistency, we also warp rendered RGB and depth to co-visible frames to construct loss with corresponding frames following similar approach in Eq.(8):
+$\mathcal{L}_{\text{BA-rgb}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},c)-\mathcal{G}(T_{j},c)|)},$
+$\mathcal{L}_{\text{BA-depth}}$ $=\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}{(|T_{i}^{j}\cdot\mathcal{G}(T_{i},d)-\mathcal{G}(T_{j},d)|)},$ (9)
+where $\mathcal{G}(T_{i},c)$ and $\mathcal{G}(T_{i},d)$ represents rendered RGB and depth of frame $i$.
+Therefore, overall loss function $\mathcal{L}_{\text{BA}}$ for joint optimization of corresponding frame poses and 3D Gaussian scene representation is the weighted sum of the above losses:
 $\mathcal{L}_{\text{BA}}=\lambda_{e}\mathcal{L}_{\text{BA-sem}}+\lambda_{c}\mathcal{L}_{\text{BA-rgb}}+\lambda_{d}\mathcal{L}_{\text{BA-depth}},$ (10)
 where $\lambda_{e}$, $\lambda_{c}$, $\lambda_{d}$ are weighting coefficients.
-This design leverages fast rendering capability of 3D Gaussian to achieve joint optimization of scene representation and camera poses in real time.
+This design leverages fast rendering capability of 3D Gaussian to achieve joint optimization of scene representation and camera poses.
 Furthermore, semantic-informed BA integrates consistency and correlation of multi-perspective semantic, geometry, and appearance information, resulting in low-drift tracking and consistent mapping.
-Table 1: Quantitative comparison of tracking accuracy for our proposed SemGauss-SLAM with other dense visual SLAM and semantic SLAM methods on Replica dataset [21].
-We utilize RMSE (cm) metric.
-To ensure more objectivity in the results, each scene is tested and averaged with three independent runs.
-Our work outperforms previous radiance field-based semantic SLAM methods across all scenes.
-Best results are highlighted as first, second.
-Methods room0 room1 room2 office0 office1 office2 office3 office4 Avg.
-Visual SLAM iMAP[23] 6.33 3.46 2.65 3.31 1.42 7.17 6.32 2.55 4.15
-NICE-SLAM [33] 1.86 2.37 2.26 1.50 1.01 1.85 5.67 3.53 2.51
-Vox-Fusion [29] 1.37 1.90 1.47 1.35 1.76 1.18 1.11 1.64 1.47
-Co-SLAM [25] 0.72 0.85 1.02 0.69 0.56 2.12 1.62 0.87 1.06
-ESLAM [6] 0.76 0.71 0.56 0.53 0.49 0.58 0.74 0.64 0.62
-Point-SLAM [19] 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72 0.52
-SplaTAM [7] 0.31 0.40 0.29 0.47 0.27 0.29 0.32 0.55 0.36
-Semantic SLAM SNI-SLAM [32] 0.50 0.55 0.45 0.35 0.41 0.33 0.62 0.50 0.46
-DNS SLAM [9] 0.49 0.46 0.38 0.34 0.35 0.39 0.62 0.60 0.45
-SemGauss-SLAM (Ours) 0.26 0.42 0.27 0.34 0.17 0.32 0.36 0.49 0.33
 4 Experiments
@@
 Datasets.
-We evaluate the performance of SemGauss-SLAM on two datasets with semantic ground truth annotations, including 8 scenes on simulated dataset Replica [21] and 5 scenes on real-world dataset ScanNet [2].
+We evaluate the performance of SemGauss-SLAM on two datasets with semantic ground truth annotations, including 8 scenes on simulated dataset Replica straub2019replica and 5 scenes on real-world dataset ScanNet dai2017scannet .
 Metrics.
-To evaluate the SLAM system and rendering quality, we follow metrics from [19].
-For mapping evaluation, we use Depth L1 (cm).
-For tracking accuracy evaluation, we utilize ATE RMSE (cm) [22].
+We follow metrics from keetha2023splatam to evaluate the SLAM system accuracy and rendering quality.
+For reconstruction metric, we use Depth L1 (cm).
+For tracking accuracy evaluation, we use ATE RMSE (cm) sturm2012benchmark .
 Moreover, we use PSNR (dB), SSIM, and LPIPS for evaluating RGB image rendering performance.
-Semantic segmentation is evaluated with respect to mIoU accuracy.
+Semantic segmentation is evaluated with respect to mIoU (%) metric.
 Baselines.
-We compare our method with the existing state of-the-art dense visual SLAM, including NeRF-based SLAM [23, 33, 25, 6, 19, 29] and 3D Gaussian SLAM [7].
-For comparison of dense semantic SLAM performance, we consider NeRF-based semantic SLAM including SNI-SLAM [32], DNS SLAM [9], and NIDS-SLAM [4] as baseline.
+We compare our method with the existing state of-the-art dense visual SLAM, including NeRF-based SLAM zhu2022nice ; wang2023co ; eslam ; sandstrom2023point ; yang2022vox and 3DGS-based SLAM keetha2023splatam .
+For comparison of dense semantic SLAM performance, we consider NeRF-based semantic SLAM zhu2023sni ; li2023dns ; haghighi2023neural , and SGS-SLAM li2024sgs which is a work parallel to ours, as baseline.
 Implementation Details.
@@
 Moreover, we set $\lambda_{e}=0.004$, $\lambda_{c}=0.5$, $\lambda_{c}=1$ in semantic-informed bundle adjustment.
-Please refer to the supplementary for further details of our implementation.
+Table 1: Quantitative comparison of SLAM accuracy and rendering performance for our proposed SemGauss-SLAM and other radiance field-based SLAM methods.
+The results are an average of 8 scenes on Replica dataset straub2019replica .
+To ensure the reliability of the results, each scene is tested and averaged with five independent runs.
+Our work outperforms previous works of all metrics.
+Methods Reconstruction Localization Rendering Quality
+Depth L1[cm] $\downarrow$ ATE RMSE[cm] $\downarrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+NICE-SLAM zhu2022nice 2.97 2.51 24.42 0.809 0.233
+Vox-Fusion yang2022vox 2.46 1.47 24.41 0.801 0.236
+Co-SLAM wang2023co 1.51 1.06 30.24 0.939 0.252
+ESLAM eslam 0.95 0.62 29.08 0.929 0.239
+SplaTAM keetha2023splatam 0.73 0.41 34.11 0.968 0.102
+SNI-SLAM zhu2023sni 0.77 0.46 29.43 0.935 0.235
+SemGauss-SLAM (Ours) 0.50 0.33 35.03 0.982 0.062
+Figure 3: Qualitative comparison on rendering quality of our method and baseline.
+We visualize 4 selected scenes of Replica dataset straub2019replica and details are highlighted with red color boxes.
+Our method achieves photo-realistic rendering quality and higher completion of reconstruction, especially in areas with rich textural information.
 4.2 Experimental Results
-Tracking Results.
+SLAM and Rendering Quality Results.
 As shown in Tab.
-1 and 2, our method achieves up to 47% relative increase in tracking accuracy compared with other dense semantic SLAM methods.
-This improvement is attributed to semantic-informed bundle adjustment, which provides multi-view constraint leveraging consistency of semantic information to reduce accumulated drift in tracking process.
-Table 2: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on ScanNet dataset [2] for tracking metric RMSE (cm).
-The results are an average of three independent runs.
-Best results are highlighted as first, second, third.
-Methods scene0000 scene0059 scene0169 scene0181 scene0207 Avg.
-NICE-SLAM [33] 12.00 14.00 10.90 13.40 6.20 11.30
-Vox-Fusion [29] 68.84 24.18 27.28 23.30 9.41 30.60
-Point-SLAM [19] 10.24 7.81 22.16 14.77 9.54 12.90
-SplaTAM [7] 12.83 10.10 12.08 11.10 7.46 10.71
-SemGauss-SLAM (Ours) 12.56 7.97 9.05 9.78 8.97 9.67
-Rendering Quality Results.
-Tab.
-3 shows rendering quality on the input views of Replica dataset [21].
-Our method achieves the best performance in PSNR, SSIM, and LPIPS compared with other dense semantic SLAM.
-Table 3: Quantitative comparison of training view rendering performance on Replica dataset [21].
-Our work outperforms other radiance field-based semantic SLAM methods on all three metrics across all scenes.
-Methods Metrics room0 room1 room2 office0 office1 office2 office3 office4 Avg.
-Visual SLAM NICE-SLAM [33] PSNR $\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
-SSIM $\uparrow$ 0.689 0.757 0.814 0.874 0.886 0.797 0.801 0.856 0.809
-LPIPS $\downarrow$ 0.330 0.271 0.208 0.229 0.181 0.235 0.209 0.198 0.233
-Vox-Fusion [29] PSNR $\uparrow$ 22.39 22.36 23.92 27.79 29.83 20.33 23.47 25.21 24.41
-SSIM $\uparrow$ 0.683 0.751 0.798 0.857 0.876 0.794 0.803 0.847 0.801
-LPIPS $\downarrow$ 0.303 0.269 0.234 0.241 0.184 0.243 0.213 0.199 0.236
-Co-SLAM [25] PSNR $\uparrow$ 27.27 28.45 29.06 34.14 34.87 28.43 28.76 30.91 30.24
-SSIM $\uparrow$ 0.910 0.909 0.932 0.961 0.969 0.938 0.941 0.955 0.939
-LPIPS $\downarrow$ 0.324 0.294 0.266 0.209 0.196 0.258 0.229 0.236 0.252
-ESLAM [6] PSNR $\uparrow$ 25.32 27.77 29.08 33.71 30.20 28.09 28.77 29.71 29.08
-SSIM $\uparrow$ 0.875 0.902 0.932 0.960 0.923 0.943 0.948 0.945 0.929
-LPIPS $\downarrow$ 0.313 0.298 0.248 0.184 0.228 0.241 0.196 0.204 0.239
-SplaTAM [7] PSNR $\uparrow$ 32.86 33.89 35.25 38.26 39.17 31.97 29.70 31.81 34.11
-SSIM $\uparrow$ 0.978 0.969 0.979 0.977 0.978 0.968 0.949 0.949 0.968
-LPIPS $\downarrow$ 0.072 0.103 0.081 0.092 0.093 0.102 0.121 0.152 0.102
-Semantic SLAM SNI-SLAM [32] PSNR $\uparrow$ 25.91 28.17 29.15 33.86 30.34 29.10 29.02 29.87 29.43
-SSIM $\uparrow$ 0.885 0.910 0.938 0.965 0.927 0.950 0.950 0.952 0.935
-LPIPS $\downarrow$ 0.307 0.292 0.245 0.182 0.225 0.238 0.192 0.198 0.235
-SemGauss-SLAM (Ours) PSNR $\uparrow$ 32.55 33.92 35.15 39.18 39.87 32.97 31.60 35.00 35.03
-SSIM $\uparrow$ 0.979 0.979 0.987 0.989 0.990 0.979 0.972 0.978 0.982
-LPIPS $\downarrow$ 0.055 0.054 0.045 0.048 0.050 0.069 0.078 0.093 0.062
-Reconstruction Results.
-As shown in Tab.
-4, our method achieves up to 77.3% increase in reconstruction accuracy compared with other radiance field-based methods.
-This enhancement in performance is due to the incorporation of semantic-informed BA process, achieving joint optimization for both camera poses and scene representation.
-Specifically, our method exploits the geometric consistency of co-visible frames to construct geometry constraints for more precise reconstruction.
-Table 4: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on Replica [21] for reconstruction metric Depth L1 (cm).
-The results are an average of three independent runs.
-Our method outperforms other methods by up to 77.3%.
-Methods room0 room1 room2 office0 office1 office2 office3 office4 Avg.
-Visual SLAM NICE-SLAM [33] 1.81 1.44 2.04 1.39 1.76 8.33 4.99 2.01 2.97
-Vox-Fusion [29] 1.09 1.90 2.21 2.32 3.40 4.19 2.96 1.61 2.46
-Co-SLAM [25] 1.05 0.85 2.37 1.24 1.48 1.86 1.66 1.54 1.51
-ESLAM [6] 0.73 0.74 1.26 0.71 1.02 0.93 1.03 1.18 0.95
-Semantic SLAM SNI-SLAM [32] 0.55 0.58 0.87 0.55 0.97 0.89 0.75 0.97 0.77
-SemGauss-SLAM (Ours) 0.54 0.46 0.43 0.29 0.22 0.51 0.98 0.56 0.50
-Semantic Segmentation Results.
-As shown in Tab.
-5, our work outperforms other dense semantic SLAM methods by up to 26% on mIoU metric and achieves 95% mIoU accuracy.
-Such enhancement attributes to the integration of semantic feature embedding into 3D Gaussian for enriched semantic representation, and semantic feature-level loss for direct guidance of semantic optimization.
-Moreover, our proposed semantic-informed BA also contributes to high semantic precision, as it leverages multiple co-visible frames to construct a globally consistent semantic map, thereby achieving high-precision semantic representation.
-Table 5: Quantitative comparison of input views semantic segmentation performance on Replica [21] for semantic metric mIoU(%).
-For one scene, we calculate mIoU between splatted semantic maps and ground truth labels, which is calculated every 4 frames, to obtain average mIoU.
-Our method achieves highest semantic accuracy across all scenes.
-Methods room0 room1 room2 office0 office1 office2 office3 office4
-NIDS-SLAM [4] 82.45 84.08 76.99 85.94
-DNS SLAM [9] 88.32 84.90 81.20 84.66
-SNI-SLAM [32] 88.42 87.43 86.16 87.63 78.63 86.49 74.01 80.22
-SemGauss-SLAM (Ours) 92.81 94.10 94.72 95.23 90.11 94.93 92.93 94.82
-Table 6: Quantitative comparison of semantic novel view synthesis performance on Replica [21] for semantic metric mIoU(%).
-For one scene, we randomly choose 100 new viewpoints for semantic novel view synthesis evaluation.
-Our method outperforms SNI-SLAM [32] across all scenes, showing excellent 3D semantic mapping accuracy.
-Methods room0 room1 room2 office0 office1 office2 office3 office4 Avg.
-SNI-SLAM [32] 51.20 50.10 54.80 70.21 63.47 58.91 64.31 71.04 60.51
-SemGauss-SLAM (Ours) 89.63 84.72 86.51 93.60 89.57 92.80 92.41 92.10 90.17
+1, our method achieves the highest accuracy in all metrics compared with other radiance field-based SLAM and up to 35% relative increase in reconstruction accuracy.
+This enhancement is attributed to the incorporation of semantic-informed bundle adjustment, which provides multi-view constraint to achieve joint optimization of
+Table 2: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on ScanNet dataset dai2017scannet for tracking metric RMSE (cm) $\downarrow$.
+Methods $0$ 0059 0169 0181 0207 Avg.
+NICE-SLAM zhu2022nice 12.00 14.00 10.90 13.40 6.20 11.30
+Vox-Fusion yang2022vox 68.84 24.18 27.28 23.30 9.41 30.60
+Point-SLAM sandstrom2023point 10.24 7.81 22.16 14.77 9.54 12.90
+SplaTAM keetha2023splatam 12.83 10.10 12.08 11.10 7.46 10.71
+SemGauss-SLAM (Ours) 11.87 7.97 8.70 9.78 8.97 9.46
+both camera poses and scene representation. Tab. 2 demonstrates our method outperforms baseline methods on real-world dataset ScanNet dai2017scannet .
 Novel View Semantic Evaluation Results.
 As shown in Tab.
-6, our method achieves up to 75% increase in semantic segmentation accuracy for semantic novel view synthesis compared with existing dense semantic SLAM.
+3, our method achieves up to 49% relative increase in semantic novel view synthesis compared with existing dense semantic SLAM.
 By introducing 3D Gaussian feature embedding for semantic representation, our method enables continuous semantic modeling.
 This modeling is crucial for generating semantically coherent scenes, as it reduces the occurrence of sharp transitions that can cause inconsistencies, ensuring semantic representation accuracy from different viewpoints for high-precision novel view semantic synthesis.
-Figure 4: Qualitative comparison on rendering quality of our method and baseline.
-We visualize 4 selected scenes of Replica dataset[21] and details are highlighted with red color boxes.
-Our method achieves photo-realistic rendering quality compared with other methods, especially in areas with rich textural information.
+Table 3: Quantitative comparison of semantic novel view synthesis performance on Replica straub2019replica for semantic metric mIoU(%).
+For one scene, we randomly choose 100 new viewpoints that are different from SLAM mapping perspectives for evaluation and mIoU is calculated between splatted semantic labels and ground truth labels.
+Our method outperforms baseline across all scenes, showing excellent 3D semantic mapping accuracy.
+Methods room0 room1 room2 office0 office1 office2 office3 office4 Avg.
+SNI-SLAM zhu2023sni 51.20 50.10 54.80 70.21 63.47 58.91 64.31 71.04 60.51
+SemGauss-SLAM (Ours) 89.63 84.72 86.51 93.60 89.57 92.80 92.41 92.10 90.17
+Table 4: Quantitative comparison of input views semantic segmentation performance on 4 scenes of Replica straub2019replica for semantic metric mIoU(%), as its results are only reported on such scenes.
+For one scene, we calculate mIoU between splatted semantic maps and ground truth labels every 4 frames to obtain average mIoU.
+Our method achieves the highest semantic accuracy across all scenes.
+Methods room0 room1 room2 office0 Avg.
+GT NIDS-SLAM haghighi2023neural 82.45 84.08 76.99 85.94 82.37
+DNS SLAM li2023dns 88.32 84.90 81.20 84.66 84.77
+SNI-SLAM zhu2023sni 88.42 87.43 86.16 87.63 87.41
+SGS-SLAM li2024sgs 92.95 92.91 92.10 92.90 92.72
+Ours (GT) 96.30 95.82 96.51 96.72 96.34
+Seg Ours 92.81 94.10 94.72 95.23 94.22
+Figure 4: Qualitative comparison on semantic novel view synthesis of our method and baseline semantic SLAM method SNI-SLAM zhu2023sni on 4 scenes of Replica straub2019replica .
+Semantic Segmentation Results.
+Since haghighi2023neural ; li2023dns ; li2024sgs ; zhu2023sni employs ground truth labels for supervision, for a fair comparison, we also present our results using ground truth labels for supervision, denoted as Our (GT).
+The results of using semantic segmentation results for supervision are Ours.
+As shown in Tab.
+4, our work outperforms existing radiance field-based semantic SLAM methods.
+Such enhancement attributes to the integration of semantic feature embedding into 3D Gaussian for enriched semantic representation, and semantic feature-level loss for direct guidance of semantic optimization.
+Moreover, our proposed semantic-informed BA also contributes to high semantic precision, as it leverages multiple co-visible frames to construct a globally consistent semantic map for high-precision semantic representation.
 Visualization.
@@
 3 shows rendering quality comparison of 4 scenes with interesting regions highlighted with colored boxes.
-For small objects, such as vases and items on the table, other methods fail to reconstruct them clearly.
-Our method leverages the high-quality rendering capability of 3D Gaussian representation and introduces semantic-informed BA to achieve detailed geometric reconstruction results.
-Specifically, our method enhances multi-view geometry consistency by BA to ensure that the reconstructed geometry aligns accurately across all observed viewpoints, leading to precise geometry reconstruction.
+For less-frequent observed areas, such as floor and side of sofa, other methods either result in low-quality reconstruction or leave holes.
+Our method leverages the high-quality rendering capability of 3D Gaussian representation and introduces semantic-informed BA to achieve detailed and complete geometric reconstruction results.
+Specifically, our method enhances multi-view geometry consistency by BA to ensure that the reconstructed geometry aligns accurately across all observed viewpoints, leading to precise and complete geometry reconstruction.
 As shown in Fig.
-5, our method achieves superior novel view semantic segmentation accuracy compared with the baseline SNI-SLAM [32].
+4, our method achieves superior novel view semantic segmentation accuracy compared with baseline SNI-SLAM zhu2023sni .
 It can be observed that SNI-SLAM struggles with segmenting ceilings in novel view synthesis as they are less frequently observed during the mapping process.
@@
 Our method introduces semantic-informed BA to construct multi-view constraints, enabling areas that are sparsely observed to effectively utilize the limited information from co-visible frames, thus establishing sufficient constraints for accurate semantic reconstruction.
-Figure 5: Qualitative comparison on semantic novel view synthesis of our method and baseline semantic SLAM method SNI-SLAM [32] on 4 scenes of Replica[21].
-Figure 6: Semantic rendering results and ground truth labels of feature-level loss ablation on two scenes of Replica [21].
 4.3 Ablation Study
-We perform ablation study on three scenes of Replica dataset[21] in Tab.
+Figure 5: Semantic rendering results and ground truth labels of feature-level loss ablation on two scenes of Replica straub2019replica .
+We perform ablation study on three scenes of Replica dataset straub2019replica in Tab.
 5 to validate the effectiveness of feature-level loss and semantic-informed BA in SemGauss-SLAM.
 Moreover, ablation of semantic-informed BA component is conducted in Tab.
-8.
+6.
 Feature-level Loss.
@@
 As shown in Fig.
-6, utilizing feature-level loss can achieve improved boundary segmentation and finer segmentation of small objects.
+5, utilizing feature-level loss can achieve improved boundary segmentation and finer segmentation of small objects.
 This enhancement occurs because feature loss compels the scene representation to capture high-dimensional and direct information within the feature space, leading to the capability of distinguishing intricate details and subtle variations within the scene.
@@
 As shown in Tab.
-8, lacking semantic constraint leads to a significant decrease in tracking performance compared with the absence of color and depth constraints.
+6, lacking semantic constraint leads to a significant decrease in tracking performance compared with the absence of color and depth constraints.
 Such result suggests that multi-view semantic constraints provide more comprehensive and accurate information, due to the consistency of semantics across multiple perspectives.
 Moreover, the absence of semantic constraint can lead to reduced semantic precision, while lacking color and depth constraints results in decreased reconstruction accuracy.
-Table 7: Ablation study of our contributions on three scenes of Replica [21].
+Table 5: Ablation study of our contributions on three scenes of Replica straub2019replica .
 Methods room0 room1 office1
@@
 SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.42 0.46 94.10 0.17 0.22 90.11
-Table 8: Ablation study of semantic-informed BA on three scenes of Replica [21].
+Table 6: Ablation study of semantic-informed BA on three scenes of Replica straub2019replica .
 (w/o semantic) without adding semantic constraint loss $\mathcal{L}_{\text{BA-sem}}$; (w/o RGB and depth) without adding RGB and depth constraint loss $\mathcal{L}_{\text{BA-rgb}}$ and $\mathcal{L}_{\text{BA-depth}}$.
@@
 We propose SemGauss-SLAM, a novel dense semantic SLAM system utilizing 3D Gaussian representation that enables dense visual mapping, robust camera tracking, and 3D semantic mapping of the whole scene.
-We incorporate semantic feature embedding into 3D Gaussian for Gaussian semantic representation, leading to real-time dense semantic mapping.
+We incorporate semantic feature embedding into 3D Gaussian for Gaussian semantic representation for dense semantic mapping.
 Moreover, we propose feature-level loss for 3D Gaussian scene optimization to achieve accurate semantic representation.

### v4

@@
 Furthermore, we propose feature-level loss for updating 3D Gaussian representation, enabling higher-level guidance for 3D Gaussian optimization.
-In addition, to reduce cumulative drift in tracking and improve semantic reconstruction accuracy, we introduce semantic-informed bundle adjustment leveraging multi-frame semantic associations for joint optimization of 3D Gaussian representation and camera poses, leading to low-drift tracking and accurate mapping.
-Our SemGauss-SLAM method demonstrates superior performance over existing radiance field-based SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in high-precision semantic segmentation and dense semantic mapping.
-1 Introduction
-Dense semantic Simultaneous Localization and Mapping (SLAM) is a fundamental challenge for robotic systems rosinol2020kimera ; mccormac2017semanticfusion and autonomous driving bao2022semantic ; lianos2018vso ; wang2022sfgan ; jiang20223d ; jiang20243dsflabelling ; jiang2022ffpa ; feng2024rogs .
+In addition, to reduce cumulative drift in tracking and improve semantic reconstruction accuracy, we introduce semantic-informed bundle adjustment.
+By leveraging multi-frame semantic associations, this strategy enables joint optimization of 3D Gaussian representation and camera poses, resulting in low-drift tracking and accurate semantic mapping.
+Our SemGauss-SLAM demonstrates superior performance over existing radiance field-based SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in high-precision semantic segmentation and dense semantic mapping.
+Code will be available at https://github.com/IRMVLab/SemGauss-SLAM.
+I Introduction
+Dense semantic Simultaneous Localization and Mapping (SLAM) is a fundamental challenge for robotic systems [1, 2] and autonomous driving [3, 4].
 It integrates semantic understanding of the environment into dense map reconstruction and performs pose estimation simultaneously.
-Traditional dense semantic SLAM has limitations including its inability to predict unknown areas liu2023unsupervised .
-Subsequent semantic SLAM based on Neural Radiance Fields (NeRF) mildenhall2021nerf methods zhu2023sni ; li2023dns ; haghighi2023neural have addressed these drawbacks, but suffer from inefficient per-pixel raycasting rendering for real-time optimization of implicit scene representation and low-quality novel view semantic representation.
-Recently, a novel radiance field based on 3D Gaussian Splatting (3DGS) kerbl20233d has demonstrated remarkable capability in scene representation, enabling high-quality and efficient rendering via splatting.
-Following the advantages of 3D Gaussian representation, 3DGS-based SLAM methods yan2023gs ; keetha2023splatam ; yugay2023gaussian ; matsuki2023gaussian ; huang2023photo have been developed to achieve photo-realistic mapping.
+Traditional dense semantic SLAM has limitations including its inability to predict unknown areas [5].
+Subsequent semantic SLAM based on Neural Radiance Fields (NeRF) [6] methods [7, 8, 9] have addressed these drawbacks, but suffer from inefficient per-pixel raycasting rendering for real-time optimization of implicit scene representation and low-quality novel view semantic representation.
+Recently, a novel radiance field based on 3D Gaussian Splatting (3DGS) [10] has demonstrated remarkable capability in scene representation, enabling high-quality and efficient rendering via splatting.
+Following the advantages of 3D Gaussian representation, 3DGS-based SLAM methods [11, 12, 13, 14, 15] have been developed to achieve photo-realistic mapping.
 However, most existing 3DGS SLAM systems focus on visual mapping to obtain RGB maps, where color information alone is insufficient for downstream tasks such as navigation.
@@
 To overcome this challenge, we introduce a novel dense semantic SLAM system based on 3D Gaussian Splatting with semantic-informed bundle adjustment, which enables high accuracy in tracking, dense mapping, and semantic segmentation from novel view.
-Different from concurrent work SGS-SLAM li2024sgs , which uses color for semantic representation to achieve 3DGS semantic SLAM, our method integrates semantic feature embedding into 3D Gaussian for semantic modeling, resulting in a more discriminative and comprehensive understanding of the scene.
+Different from concurrent work SGS-SLAM [16], which uses color for semantic representation to achieve 3DGS semantic SLAM, our method integrates semantic feature embedding into 3D Gaussian for semantic modeling, resulting in a more discriminative and comprehensive understanding of the scene.
 Moreover, by leveraging the explicit 3D structure of 3D Gaussian representation, our embedding-based semantic Gaussian representation is capable of capturing the spatial distribution of semantic information, achieving high-accuracy novel view semantic segmentation.
@@
   • We perform semantic-informed bundle adjustment by leveraging multi-view semantic constraints for joint optimization of camera poses and 3D Gaussian representation, achieving low-drift tracking and accurate semantic mapping.
-  • We conduct extensive evaluations on two challenging datasets, Replica straub2019replica and ScanNet dai2017scannet , to demonstrate our method achieves state-of-the-art performance compared with existing radiance field-based SLAM in mapping, tracking, semantic segmentation and novel-view synthesis.
-2 Related Work
+  • We conduct extensive evaluations on challenging datasets, to demonstrate our method achieves superior performance compared with existing radiance field-based SLAM in mapping, tracking, semantic segmentation, and novel-view synthesis.
+II Related Works
 Traditional Semantic SLAM.
-Traditional semantic SLAM employs explicit 3D representation such as surfels mccormac2017semanticfusion , mesh rosinol2020kimera ; tian2022kimera ; hughes2022hydra and Truncated Signed Distance Fields (TSDF) schmid2022panoptic ; mccormac2018fusion++ ; grinvald2019volumetric for dense semantic mapping.
-SemanticFusion mccormac2017semanticfusion uses surfel representation and employs Conditional Random Field (CRF) for updating class probability distribution incrementally.
-Fusion++ mccormac2018fusion++ performs object-level SLAM where each object is reconstructed within its own TSDF volume and segmented based on estimated foreground probability.
-Kimera rosinol2020kimera utilizes visual-inertial odometry for pose estimation and generates dense semantic mesh maps.
+Traditional semantic SLAM employs explicit 3D representation such as surfels [2], mesh [1, 17, 18] and Truncated Signed Distance Fields (TSDF) [19, 20, 21] for dense semantic mapping.
+SemanticFusion [2] uses surfel representation and employs Conditional Random Field (CRF) for updating class probability distribution incrementally.
+Fusion++ [20] performs object-level SLAM where each object is reconstructed within its own TSDF volume and segmented based on estimated foreground probability.
+Kimera [1] utilizes visual-inertial odometry for pose estimation and generates dense semantic mesh maps.
 However, these explicit scene modeling methods not only require high storage space but also fail to achieve high-fidelity and complete 3D reconstruction.
 Neural Implicit SLAM.
-Neural implicit representation mildenhall2021nerf has shown promising capability in scene reconstruction for dense mapping.
-iMAP imap first achieves real-time mapping and tracking utilizing a single MLP network for scene representation.
-To overcome over-smoothed scene reconstruction and improve scalability, NICE-SLAM zhu2022nice adopts hierarchical feature grid representation.
-Following this, several works yang2022vox ; eslam ; wang2023co ; sandstrom2023point ; kong2023vmap ; hu2024cp ; deng2023plgslam introduce more efficient scene representation, such as hash-based feature grid and feature plane, to achieve more accurate SLAM performance.
-For neural implicit semantic SLAM, DNS SLAM li2023dns utilizes 2D semantic priors and integrates multi-view geometry constraints for semantic reconstruction.
-SNI-SLAM zhu2023sni introduces feature collaboration and one-way correlation decoder for improved scene representation in semantic mapping.
+Neural implicit representation [6] has shown promising capability in scene reconstruction for dense mapping.
+iMAP [22] first achieves mapping and tracking utilizing a single MLP network for scene representation.
+To overcome over-smoothed scene reconstruction and improve scalability, NICE-SLAM [23] adopts hierarchical feature grid representation.
+Following this, several works [24, 25, 26, 27, 28, 29] introduce more efficient scene representation, such as hash-based feature grid and feature plane, to achieve more accurate SLAM performance.
+For neural implicit semantic SLAM, DNS SLAM [8] utilizes 2D semantic priors and integrates multi-view geometry constraints for semantic reconstruction.
+SNI-SLAM [7] introduces feature collaboration and one-way correlation decoder for improved scene representation in semantic mapping.
 However, these methods require specific scene bounds for mapping and suffer from cumulative drift in pose estimation.
@@
 3D Gaussian Splatting SLAM.
-3D Gaussian Splatting kerbl20233d emerges as a promising 3D scene representation using a set of 3D Gaussians with learnable properties, including 3D center position, anisotropic covariance, opacity, and color.
-This representation is capable of quick differential rendering through splatting and has a wide range of applications in dynamic scene modeling luiten2023dynamic ; wu20234d ; yang2023deformable ; yan2024street ; zhou2023drivinggaussian , semantic segmentation cen2023segment ; ye2023gaussian ; zhou2023feature , scene editing zhuang2024tip ; fang2023gaussianeditor ; huang2023point .
-Our main focus is on 3D Gaussian SLAM yan2023gs ; keetha2023splatam ; matsuki2023gaussian ; yugay2023gaussian .
+3D Gaussian Splatting [10] emerges as a promising 3D scene representation using a set of 3D Gaussians with learnable properties, including 3D center position, anisotropic covariance, opacity, and color.
+This representation is capable of quick differential rendering through splatting and has a wide range of applications in dynamic scene modeling [30, 31], semantic segmentation [32, 33] and scene editing [34].
+Our main focus is on 3D Gaussian SLAM [11, 12, 14, 13].
 These works emerge concurrently and all perform dense visual SLAM by leveraging scene geometry and appearance modeling capabilities of 3D Gaussian representation.
-SplaTAM keetha2023splatam introduces silhouette-guided optimization to facilitate structured map expansion for dense mapping of visual SLAM.
-Gaussian Splatting SLAM matsuki2023gaussian performs novel Gaussian insertion and pruning for monocular SLAM.
+SplaTAM [12] introduces silhouette-guided optimization to facilitate structured map expansion for dense mapping of visual SLAM.
+Gaussian Splatting SLAM [14] performs novel Gaussian insertion and pruning for monocular SLAM.
 However, the capability of 3D Gaussian representation extends well beyond appearance and geometry modeling of the scene, as it can be augmented to perform semantic scene understanding.
-The concurrent work SGS-SLAM li2024sgs introduces semantic color associated with the Gaussian for semantic representation.
+The concurrent work SGS-SLAM [16] introduces semantic color associated with the Gaussian for semantic representation.
 However, this color-based semantic modeling method ignores the higher-level information inherent in semantics, which is insufficient for semantic representation of the environment.
@@
 During the SLAM process, we utilize semantic-informed bundle adjustment based on multi-frame constraints for joint optimization of poses and 3D Gaussian representation.
-3 Method
+III Method
 The overview of SemGauss-SLAM is shown in Fig.
@@
 Sec.
-3.1 introduces semantic Gaussian representation for dense semantic mapping, as well as the process of 3DGS-based semantic SLAM.
+III-A introduces semantic Gaussian representation for dense semantic mapping, as well as the process of 3DGS-based semantic SLAM.
 Sec.
-3.2 introduces loss functions in the mapping and tracking process.
+III-B introduces loss functions in the mapping and tracking process.
 Sec.
-3.3 presents semantic-informed bundle adjustment to reduce accumulated drift in tracking and improve mapping accuracy.
-3.1 3D Gaussian Semantic Mapping and Tracking
+III-C presents semantic-informed bundle adjustment to reduce accumulated drift in tracking and improve mapping accuracy.
+III-A 3D Gaussian Semantic Mapping and Tracking
 Semantic Gaussian Representation.
@@
 Therefore, instead of initializing these feature embeddings randomly, we propagate 2D semantic features extracted from images to 3D Gaussian as the initial values to achieve faster convergence of semantic Gaussian optimization.
-In this work, we use a universal feature extractor Dinov2 oquab2023dinov2 , followed by a pretrained classifier to construct the segmentation network.
+In this work, we use a universal feature extractor DINOv2 [35], followed by a pretrained classifier to construct the segmentation network.
 Overall, each Gaussian includes 3D center position $\mu$, radius $r$, color $c=(r,g,b)$, opacity $\alpha$, 16-channel semantic feature embedding $e$, and is defined as standard Gaussian equation multiplied by opacity $\alpha$:
@@
 3D Gaussian Rendering.
-Following Gaussian splatting kerbl20233d , we project 3D Gaussians to 2D splats for high-fidelity differentiable rendering, which facilitates explicit gradient flow for Gaussian scene optimization and pose estimation.
+Following Gaussian splatting [10], we project 3D Gaussians to 2D splats for high-fidelity differentiable rendering, which facilitates explicit gradient flow for Gaussian scene optimization and pose estimation.
 Specifically, for splatting of semantic feature embedding, we first sort all Gaussians in order of depth from near to far and then blend $N$ ordered points projecting to pixel $p=(u,v)$, obtaining 2D semantic feature maps $E(p)$:
@@
 For RGB and depth rendering, we follow the similar approach in Eq.(2):
-$C(p)=\sum_{i\in N}c_{i}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)),\ \ D(p)=\sum_{i\in N}d_{i}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)),$ (4)
+$C(p)=\sum_{i\in N}c_{i}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)),$ (4)
+$D(p)=\sum_{i\in N}d_{i}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p))$
 where $C(p)$ and $D(p)$ represents splatted 2D color and depth images respectively.
 Moreover, given the camera pose, visibility information of Gaussian is required for mapping and tracking process, such as adding new Gaussian and loss construction.
-Therefore, following  keetha2023splatam , we render a silhouette image to determine visibility:
+Therefore, following  [12], we render a silhouette image to determine visibility:
 $Sil(p)=\sum_{i\in N}g_{i}(p)\prod_{j=1}^{i-1}(1-g_{j}(p)).$ (5)
 Tracking Process.
-During tracking, we keep 3D Gaussian parameters fixed and only optimize camera pose $T$ of current frame.
-Since the proximity between adjacent frames is relatively small, we assume a constant velocity model to obtain an initial pose estimation for the new frame.
-The camera pose is then iteratively refined by optimization through loss construction between differentiable rendering from 3D Gaussian and camera observation within the visible silhouette.
+During tracking process, we keep 3D Gaussian parameters fixed and only optimize camera pose of current frame.
+Since the proximity between adjacent frames is relatively small, we assume a constant velocity model to obtain an initial pose estimation for the current frame.
+The camera pose is then iteratively refined through optimization by minimizing the loss between two components: differentiable rendering from 3D Gaussian representation and camera observation within the visible silhouette.
 Mapping Process.
@@
 Then, when the overlap of coming frame with rendering of the existing map is less than half, we add a new Gaussian for incremental mapping.
-3.2 Loss Functions
+III-B Loss Functions
 For optimization of semantic scene representation, we utilize cross-entropy loss for constructing semantic loss $\mathcal{L}_{s}$.
-Furthermore, instead of using semantic loss solely, we introduce a feature-level loss $\mathcal{L}_{f}$ for higher-level semantic optimization guidance.
-Feature-level loss is obtained by constructing L1 loss between extracted features generated by Dinov2 oquab2023dinov2 -based feature extractor and splatted features obtained from 3D Gaussian representation.
-Compared with semantic loss, feature loss provides direct guidance on intermediate features, leading to more robust and accurate semantic scene understanding.
-We employ RGB loss $\mathcal{L}_{c}$ and depth loss $\mathcal{L}_{d}$ for optimization of scene color and geometry representation.
+Furthermore, instead of solely using semantic loss for optimization, we introduce a feature-level loss $\mathcal{L}_{f}$ for higher-level guidance of semantic optimization:
+$\mathcal{L}_{f}=\sum_{p\in P_{M}}|F_{e}-E(p)|,$ (6)
+where $F_{e}$ is extracted features generated by DINOv2 [35]-based feature extractor.
+$P_{M}$ represents a set of all pixels in the rendered image.
+Compared with semantic loss, feature loss offers explicit supervision on intermediate features for semantic Gaussian optimization, leading to more robust and accurate semantic understanding of the scene.
+We employ RGB loss $\mathcal{L}_{c}$ and depth loss $\mathcal{L}_{d}$ for optimizing scene color and geometry representation.
 $\mathcal{L}_{c}$ and $\mathcal{L}_{d}$ are both L1 loss constructed by comparing the RGB and depth splats with the input RGB-D frame.
 These loss functions are then utilized for mapping and pose estimation.
-In mapping process, we construct loss over all rendered pixels for 3D Gaussian scene optimization.
-Moreover, we add SSIM term to RGB loss following kerbl20233d .
+In the mapping process, we construct loss over all rendered pixels for 3D Gaussian optimization.
+Moreover, we add SSIM term to RGB loss following [10].
 The complete loss function for mapping is a weighted sum of the above losses:
-$\mathcal{L}_{\text{mapping}}=\sum_{p\in P_{M}}{(\lambda_{f_{m}}\mathcal{L}_{f}(p)+\lambda_{s_{m}}\mathcal{L}_{s}(p)+\lambda_{c_{m}}\mathcal{L}_{c}(p)+\lambda_{d_{m}}\mathcal{L}_{d}(p))}.$ (6)
-$P_{M}$ represents a set of all pixels in rendered image. $\lambda_{f_{m}}$, $\lambda_{s_{m}}$, $\lambda_{c_{m}}$, $\lambda_{d_{m}}$ are weighting coefficients.
-During tracking, utilizing an overly constrained loss function for pose estimation can lead to decreased accuracy in camera pose and increased processing time.
+$\footnotesize\mathcal{L}_{\text{mapping}}=\sum_{p\in P_{M}}{(\lambda_{f_{m}}\mathcal{L}_{f}(p)+\lambda_{s_{m}}\mathcal{L}_{s}(p)+\lambda_{c_{m}}\mathcal{L}_{c}(p)+\lambda_{d_{m}}\mathcal{L}_{d}(p))}.$ (7)
+$\lambda_{f_{m}}$, $\lambda_{s_{m}}$, $\lambda_{c_{m}}$, $\lambda_{d_{m}}$ are weighting coefficients.
+In the tracking process, using an overly constrained loss function for pose estimation can lead to decreased accuracy in camera pose and increased processing time.
 Therefore, loss function for tracking is constructed based only on a weighted sum of RGB loss and depth loss:
@@
 $P_{T}$ represents pixels that are rendered from well-optimized part of 3D Gaussian map, which is area that rendered visibility silhouette $Sil(p)$ is greater than 0.99.
-3.3 Semantic-informed Bundle Adjustment
+III-C Semantic-informed Bundle Adjustment
 Currently, existing radiance field-based semantic SLAM systems utilize the latest input RGB-D frame to construct RGB and depth loss for pose estimation.
@@
 where $\lambda_{e}$, $\lambda_{c}$, $\lambda_{d}$ are weighting coefficients.
-This design leverages fast rendering capability of 3D Gaussian to achieve joint optimization of scene representation and camera poses.
+This design leverages the fast rendering capability of 3D Gaussian to achieve joint optimization of scene representation and camera poses.
 Furthermore, semantic-informed BA integrates consistency and correlation of multi-perspective semantic, geometry, and appearance information, resulting in low-drift tracking and consistent mapping.
-4 Experiments
-4.1 Experimental Setup
+IV Experiments
+IV-A Experimental Setup
 Datasets.
-We evaluate the performance of SemGauss-SLAM on two datasets with semantic ground truth annotations, including 8 scenes on simulated dataset Replica straub2019replica and 5 scenes on real-world dataset ScanNet dai2017scannet .
+We evaluate the performance of SemGauss-SLAM on two datasets with semantic ground truth annotations, including 8 scenes on simulated dataset Replica [36] and 5 scenes on real-world dataset ScanNet [37].
 Metrics.
-We follow metrics from keetha2023splatam to evaluate the SLAM system accuracy and rendering quality.
+We follow metrics from [12] to evaluate the SLAM system accuracy and rendering quality.
 For reconstruction metric, we use Depth L1 (cm).
-For tracking accuracy evaluation, we use ATE RMSE (cm) sturm2012benchmark .
+For tracking accuracy evaluation, we use ATE RMSE (cm) [38].
 Moreover, we use PSNR (dB), SSIM, and LPIPS for evaluating RGB image rendering performance.
 Semantic segmentation is evaluated with respect to mIoU (%) metric.
+Figure 3: Qualitative comparison on rendering quality of our method and baseline.
+We visualize 5 selected scenes of Replica  [36] and ScanNet [37] dataset.
+Details are highlighted with red color boxes.
+Our method achieves photo-realistic rendering quality and higher completion of reconstruction, especially in areas with rich textural information.
+Figure 4: Qualitative comparison on semantic novel view synthesis on 3 scenes of Replica [36].
 Baselines.
-We compare our method with the existing state of-the-art dense visual SLAM, including NeRF-based SLAM zhu2022nice ; wang2023co ; eslam ; sandstrom2023point ; yang2022vox and 3DGS-based SLAM keetha2023splatam .
-For comparison of dense semantic SLAM performance, we consider NeRF-based semantic SLAM zhu2023sni ; li2023dns ; haghighi2023neural , and SGS-SLAM li2024sgs which is a work parallel to ours, as baseline.
+We compare our method with the existing state of-the-art dense visual SLAM, including NeRF-based SLAM [23, 26, 25, 27, 24] and 3DGS-based SLAM [12].
+For comparison of dense semantic SLAM performance, we consider NeRF-based semantic SLAM [7, 8, 9], and SGS-SLAM [16] which is a work parallel to ours, as baseline.
 Implementation Details.
@@
 For experimental settings, we perform mapping every 8 frames.
-The weighting coefficients of each loss are $\lambda_{f_{m}}=0.01$, $\lambda_{s_{m}}=0.01$, $\lambda_{c_{m}}=0.5$, $\lambda_{d_{m}}=1$ in mapping, $\lambda_{c_{t}}=0.5$ and $\lambda_{d_{t}}=1$ in tracking.
+Weighting coefficients of each loss are $\lambda_{f_{m}}=0.01$, $\lambda_{s_{m}}=0.01$, $\lambda_{c_{m}}=0.5$, $\lambda_{d_{m}}=1$ in mapping, $\lambda_{c_{t}}=0.5$ and $\lambda_{d_{t}}=1$ in tracking.
 Moreover, we set $\lambda_{e}=0.004$, $\lambda_{c}=0.5$, $\lambda_{c}=1$ in semantic-informed bundle adjustment.
-Table 1: Quantitative comparison of SLAM accuracy and rendering performance for our proposed SemGauss-SLAM and other radiance field-based SLAM methods.
-The results are an average of 8 scenes on Replica dataset straub2019replica .
-To ensure the reliability of the results, each scene is tested and averaged with five independent runs.
-Our work outperforms previous works of all metrics.
+The above weighting coefficients are set based on the numerical scale of different losses.
+TABLE I: Comparison of SLAM accuracy and rendering performance. Results are an average of 8 scenes on Replica dataset [36]
 Methods Reconstruction Localization Rendering Quality
-Depth L1[cm] $\downarrow$ ATE RMSE[cm] $\downarrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-NICE-SLAM zhu2022nice 2.97 2.51 24.42 0.809 0.233
-Vox-Fusion yang2022vox 2.46 1.47 24.41 0.801 0.236
-Co-SLAM wang2023co 1.51 1.06 30.24 0.939 0.252
-ESLAM eslam 0.95 0.62 29.08 0.929 0.239
-SplaTAM keetha2023splatam 0.73 0.41 34.11 0.968 0.102
-SNI-SLAM zhu2023sni 0.77 0.46 29.43 0.935 0.235
+[…]
+NICE-SLAM [23] 2.97 2.51 24.42 0.809 0.233
+Vox-Fusion [24] 2.46 1.47 24.41 0.801 0.236
+Co-SLAM [26] 1.51 1.06 30.24 0.939 0.252
+ESLAM [25] 0.95 0.62 29.08 0.929 0.239
+SplaTAM [12] 0.73 0.41 34.11 0.968 0.102
+SNI-SLAM [7] 0.77 0.46 29.43 0.935 0.235
 SemGauss-SLAM (Ours) 0.50 0.33 35.03 0.982 0.062
-Figure 3: Qualitative comparison on rendering quality of our method and baseline.
-We visualize 4 selected scenes of Replica dataset straub2019replica and details are highlighted with red color boxes.
-Our method achieves photo-realistic rendering quality and higher completion of reconstruction, especially in areas with rich textural information.
-4.2 Experimental Results
+IV-B Experimental Results
 SLAM and Rendering Quality Results.
 As shown in Tab.
-1, our method achieves the highest accuracy in all metrics compared with other radiance field-based SLAM and up to 35% relative increase in reconstruction accuracy.
-This enhancement is attributed to the incorporation of semantic-informed bundle adjustment, which provides multi-view constraint to achieve joint optimization of
-Table 2: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on ScanNet dataset dai2017scannet for tracking metric RMSE (cm) $\downarrow$.
+I, our method achieves the highest accuracy in all metrics compared with other radiance field-based SLAM and up to 35% relative increase in reconstruction accuracy.
+This enhancement is attributed to the incorporation of semantic-informed bundle adjustment, which provides multi-view constraint to achieve joint optimization of both camera poses and scene representation.
+Tab.
+II demonstrates our method outperforms baseline methods on real-world dataset ScanNet [37].
+Note that each scene is tested and averaged with five independent runs to ensure the reliability of results.
+TABLE II: Accuracy comparison on ScanNet dataset [37] for tracking metric RMSE (cm) $\downarrow$.
 Methods $0$ 0059 0169 0181 0207 Avg.
-NICE-SLAM zhu2022nice 12.00 14.00 10.90 13.40 6.20 11.30
-Vox-Fusion yang2022vox 68.84 24.18 27.28 23.30 9.41 30.60
-Point-SLAM sandstrom2023point 10.24 7.81 22.16 14.77 9.54 12.90
-SplaTAM keetha2023splatam 12.83 10.10 12.08 11.10 7.46 10.71
+NICE-SLAM [23] 12.00 14.00 10.90 13.40 6.20 11.30
+Vox-Fusion [24] 68.84 24.18 27.28 23.30 9.41 30.60
+Point-SLAM [27] 10.24 7.81 22.16 14.77 9.54 12.90
+SplaTAM [12] 12.83 10.10 12.08 11.10 7.46 10.71
 SemGauss-SLAM (Ours) 11.87 7.97 8.70 9.78 8.97 9.46
-both camera poses and scene representation. Tab. 2 demonstrates our method outperforms baseline methods on real-world dataset ScanNet dai2017scannet .
 Novel View Semantic Evaluation Results.
-As shown in Tab.
-3, our method achieves up to 49% relative increase in semantic novel view synthesis compared with existing dense semantic SLAM.
-By introducing 3D Gaussian feature embedding for semantic representation, our method enables continuous semantic modeling.
-This modeling is crucial for generating semantically coherent scenes, as it reduces the occurrence of sharp transitions that can cause inconsistencies, ensuring semantic representation accuracy from different viewpoints for high-precision novel view semantic synthesis.
-Table 3: Quantitative comparison of semantic novel view synthesis performance on Replica straub2019replica for semantic metric mIoU(%).
-For one scene, we randomly choose 100 new viewpoints that are different from SLAM mapping perspectives for evaluation and mIoU is calculated between splatted semantic labels and ground truth labels.
-Our method outperforms baseline across all scenes, showing excellent 3D semantic mapping accuracy.
-Methods room0 room1 room2 office0 office1 office2 office3 office4 Avg.
-SNI-SLAM zhu2023sni 51.20 50.10 54.80 70.21 63.47 58.91 64.31 71.04 60.51
-SemGauss-SLAM (Ours) 89.63 84.72 86.51 93.60 89.57 92.80 92.41 92.10 90.17
-Table 4: Quantitative comparison of input views semantic segmentation performance on 4 scenes of Replica straub2019replica for semantic metric mIoU(%), as its results are only reported on such scenes.
-For one scene, we calculate mIoU between splatted semantic maps and ground truth labels every 4 frames to obtain average mIoU.
-Our method achieves the highest semantic accuracy across all scenes.
+Tab.
+III shows that our method achieves up to 49% relative increase in semantic novel view synthesis compared with SNI-SLAM [7], showing excellent 3D semantic mapping accuracy.
+Note that for one scene, we randomly choose 100 new viewpoints that are different from SLAM mapping perspectives for evaluation and mIoU is calculated between splatted semantic labels and ground truth labels.
+By introducing Gaussian feature embedding for semantic representation, our method enables continuous semantic modeling.
+This modeling is crucial for generating semantically coherent scenes, as it reduces the occurrence of sharp transitions that can cause inconsistencies, ensuring accurate semantic representation from novel viewpoints.
+TABLE III: Quantitative comparison of semantic novel view synthesis performance on Replica [36] for semantic metric mIoU(%).
 Methods room0 room1 room2 office0 Avg.
-GT NIDS-SLAM haghighi2023neural 82.45 84.08 76.99 85.94 82.37
-DNS SLAM li2023dns 88.32 84.90 81.20 84.66 84.77
-SNI-SLAM zhu2023sni 88.42 87.43 86.16 87.63 87.41
-SGS-SLAM li2024sgs 92.95 92.91 92.10 92.90 92.72
+SNI-SLAM [7] 51.20 50.10 54.80 70.21 56.58
+SemGauss-SLAM (Ours) 89.63 84.72 86.51 93.60 88.62
+TABLE IV: Input views semantic segmentation performance on 4 scenes of Replica [36] for semantic metric mIoU(%).
+Methods room0 room1 room2 office0 Avg.
+GT NIDS-SLAM [9] 82.45 84.08 76.99 85.94 82.37
+DNS SLAM [8] 88.32 84.90 81.20 84.66 84.77
+SNI-SLAM [7] 88.42 87.43 86.16 87.63 87.41
+SGS-SLAM [16] 92.95 92.91 92.10 92.90 92.72
 Ours (GT) 96.30 95.82 96.51 96.72 96.34
 Seg Ours 92.81 94.10 94.72 95.23 94.22
-Figure 4: Qualitative comparison on semantic novel view synthesis of our method and baseline semantic SLAM method SNI-SLAM zhu2023sni on 4 scenes of Replica straub2019replica .
 Semantic Segmentation Results.
-Since haghighi2023neural ; li2023dns ; li2024sgs ; zhu2023sni employs ground truth labels for supervision, for a fair comparison, we also present our results using ground truth labels for supervision, denoted as Our (GT).
+Since [9, 8, 16, 7] employs ground truth labels for supervision, for a fair comparison, we also present our results using ground truth labels for supervision, denoted as Our (GT).
 The results of using semantic segmentation results for supervision are Ours.
 As shown in Tab.
-4, our work outperforms existing radiance field-based semantic SLAM methods.
+IV, our work outperforms existing radiance field-based semantic SLAM methods.
 Such enhancement attributes to the integration of semantic feature embedding into 3D Gaussian for enriched semantic representation, and semantic feature-level loss for direct guidance of semantic optimization.
@@
 As shown in Fig.
-4, our method achieves superior novel view semantic segmentation accuracy compared with baseline SNI-SLAM zhu2023sni .
+4, our method achieves superior novel view semantic segmentation accuracy compared with baseline SNI-SLAM [7].
 It can be observed that SNI-SLAM struggles with segmenting ceilings in novel view synthesis as they are less frequently observed during the mapping process.
@@
 Our method introduces semantic-informed BA to construct multi-view constraints, enabling areas that are sparsely observed to effectively utilize the limited information from co-visible frames, thus establishing sufficient constraints for accurate semantic reconstruction.
-4.3 Ablation Study
-Figure 5: Semantic rendering results and ground truth labels of feature-level loss ablation on two scenes of Replica straub2019replica .
-We perform ablation study on three scenes of Replica dataset straub2019replica in Tab.
-5 to validate the effectiveness of feature-level loss and semantic-informed BA in SemGauss-SLAM.
+Figure 5: Semantic rendering results and ground truth labels of feature-level loss ablation on two scenes of Replica [36].
+TABLE V: Ablation study of our contributions on Replica [36].
+Methods room0 office1
+RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
+w/o feature-level loss 0.26 0.54 83.60 0.17 0.22 80.13
+w/o semantic-informed BA 0.35 0.70 90.10 0.21 0.30 87.73
+SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.17 0.22 90.11
+IV-C Ablation Study
+We perform ablation study on two scenes of Replica dataset [36] in Tab.
+V to validate the effectiveness of feature-level loss and semantic-informed BA in SemGauss-SLAM.
 Moreover, ablation of semantic-informed BA component is conducted in Tab.
-6.
+VI.
 Feature-level Loss.
 Tab.
-5 shows that incorporating feature-level loss significantly enhances semantic segmentation performance, while having little effect on tracking and geometric reconstruction.
+V shows that incorporating feature-level loss significantly enhances semantic segmentation performance, while having little effect on tracking and geometric reconstruction.
 This result occurs because feature-level loss influences only the optimization of semantic features, without affecting the optimization of geometry and pose estimation.
@@
 Tab.
-5 shows that introducing semantic-informed BA leads to enhancements in tracking, reconstruction, and semantic segmentation.
+V shows that introducing semantic-informed BA leads to enhancements in tracking, reconstruction, and semantic segmentation.
 This improvement is due to the joint optimization of camera poses and scene representation, which is informed by multi-view constraints.
 As shown in Tab.
-6, lacking semantic constraint leads to a significant decrease in tracking performance compared with the absence of color and depth constraints.
+VI, lacking semantic constraint leads to a significant decrease in tracking performance compared with the absence of color and depth constraints.
 Such result suggests that multi-view semantic constraints provide more comprehensive and accurate information, due to the consistency of semantics across multiple perspectives.
 Moreover, the absence of semantic constraint can lead to reduced semantic precision, while lacking color and depth constraints results in decreased reconstruction accuracy.
-Table 5: Ablation study of our contributions on three scenes of Replica straub2019replica .
-Methods room0 room1 office1
-RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
-w/o feature-level loss 0.26 0.54 83.60 0.42 0.46 86.10 0.17 0.22 80.13
-w/o semantic-informed BA 0.35 0.70 90.10 0.51 0.60 91.08 0.21 0.30 87.73
-SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.42 0.46 94.10 0.17 0.22 90.11
-Table 6: Ablation study of semantic-informed BA on three scenes of Replica straub2019replica .
-(w/o semantic) without adding semantic constraint loss $\mathcal{L}_{\text{BA-sem}}$; (w/o RGB and depth) without adding RGB and depth constraint loss $\mathcal{L}_{\text{BA-rgb}}$ and $\mathcal{L}_{\text{BA-depth}}$.
-Methods room0 room1 office1
-RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
-w/o semantic 0.35 0.66 90.15 0.50 0.49 91.50 0.20 0.24 87.95
-w/o RGB and depth 0.31 0.70 92.01 0.48 0.59 93.90 0.18 0.28 89.60
-SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.42 0.46 94.10 0.17 0.22 90.11
-5 Conclusion
+TABLE VI: Ablation study of semantic-informed BA on Replica [36].
+(w/o semantic) without adding $\mathcal{L}_{\text{BA-sem}}$; (w/o RGB and depth) without adding $\mathcal{L}_{\text{BA-rgb}}$ and $\mathcal{L}_{\text{BA-depth}}$.
+Methods room0 office1
+RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
+w/o semantic 0.35 0.66 90.15 0.20 0.24 87.95
+w/o RGB and depth 0.31 0.70 92.01 0.18 0.28 89.60
+SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.17 0.22 90.11
+V Conclusion
 We propose SemGauss-SLAM, a novel dense semantic SLAM system utilizing 3D Gaussian representation that enables dense visual mapping, robust camera tracking, and 3D semantic mapping of the whole scene.